### v1

# Emulating Full Client Participation: A Long-Term Client Selection Strategy for Federated Learning

Abstract
Client selection significantly affects the system convergence efficiency and is a crucial problem in federated learning. Existing methods often select clients by evaluating each round individually and overlook the necessity for long-term optimization, resulting in suboptimal performance and potential fairness issues. In this study, we propose a novel client selection strategy designed to emulate the performance achieved with full client participation. In a single round, we select clients by minimizing the gradient-space estimation error between the client subset and the full client set. In multi-round selection, we introduce a novel individual fairness constraint, which ensures that clients with similar data distributions have similar frequencies of being selected. This constraint guides the client selection process from a long-term perspective. We employ Lyapunov optimization and submodular functions to efficiently identify the optimal subset of clients, and provide a theoretical analysis of the convergence ability. Experiments demonstrate that the proposed strategy significantly improves both accuracy and fairness compared to previous methods while also exhibiting efficiency by incurring minimal time overhead.
Index Terms:
Federated Learning, Client Selection, Individual Fairness, Coreset Selection, Lyapunov Optimization
I Introduction
Federated learning (FL) facilitates collaborative model training without the necessity of sharing local data [1, 2], and is widely used in various domains [3, 4, 5]. In FL, model parameters or gradient updates are frequently exchanged between the central server and participating clients, leading to substantial communication overhead. To address the challenge posed by the limited bandwidth, a common approach is to select a subset of clients for local training, which reduces the communication costs compared with involving all clients [6, 7]. However, in heterogeneous scenarios, different selection strategies lead to distinct optimization trajectories, and an improper client selection strategy can cause the optimization trajectory to deviate seriously from the optimal one, significantly hindering the model performance. Therefore, a critical challenge is to develop an effective client selection strategy that reduces the communication overhead while ensuring model convergence.
In most cases, the global models generated by full client participation are considered the gold standard or benchmark. This is because they represent models trained on the entire dataset available across all clients, offering a comprehensive reference point for evaluating model performance. Building upon this understanding, we propose an intuitive strategy where clients are selected in a manner that ensures the optimization trajectory closely resembles that achieved with full client participation over the long term. We refer to this strategy as the long-term client selection strategy. This strategy yields two key advantages. First, by emulating the optimization trajectory of full client participation, it aims to achieve comparable performance. Moreover, it reduces communication overhead by involving only a subset of clients in the process.
Our Goal. In this work, we investigate the three questions: 1) what principles should be followed to implement this strategy; 2) how to mathematically formulate this strategy and provide a fast solution; and 3) is this strategy more effective than existing methods? To answer the first question, we propose two crucial principles to be followed.
Principle I: In each round, the data distribution generated by the selected subset of clients should closely resemble the data distribution of the full client participation. For example, in Fig. 1 (a), clients collaboratively train an image classification model, where each client has images of only one kind of animal, and three clusters exist in their data distributions. Due to limited bandwidth, the server is required only to select three clients in each round. In this case, as shown in Fig. 1 (b), the ideal strategy is to select one client from each cluster instead of three clients from the same cluster. This ensures that the data distribution resulting from the subset of selected clients serves as a reliable estimation of that from the full set of clients. However, in real-world applications, the clustering pattern in data distributions may be unknown or non-existent. Therefore, the key challenge here is to determine the subset of clients that best approximates the full client set without imposing extensive assumptions on their data distributions.
Figure 1: An example of the long-term client selection strategy.
However, if only Principle I is followed and client selection is performed by evaluating each round individually, there is a risk that the same subset of clients may be consistently selected across multiple rounds, while others may have minimal opportunities for selection. This phenomenon is indeed observed in our experimental results. In such cases, the system may accurately predict outcomes for frequently selected clients but makes inaccurate or biased predictions for those rarely selected, resulting in degradation of the global model’s performance and ineffective convergence. Furthermore, these seldom-selected clients may perceive unfair treatment and lose motivation to participate, ultimately endangering the sustainability of the federated system.
To mitigate this problem, we propose Principle II, which guides client selection across multiple rounds from a long-term perspective. Principle II: In multi-round selection, clients with similar data distributions should have similar frequencies of being selected. For example, as illustrated in Fig. 1, clients 1, 2, and 3 should share similar selection frequencies, as should clients 4 and 5, while the frequencies for clients 1 and 4 may differ. This principle aligns with the concept of individual fairness (IF) in central machine learning [8], which posits that similar instances or individuals in a dataset should yield similar predictions or outcomes. In our work, we refer to this principle as the individual fairness constraint. Here, the key challenge is to quantify the similarity between clients in terms of data distribution and selection frequency, respectively, and further formulate individual fairness mathematically.
This proposed individual fairness offers two key benefits. First, as illustrated in Fig. 1 (c), it promotes the diversity of selected clients across multiple rounds and reduces model prediction bias. Furthermore, existing fair federated learning approaches often utilize a uniform sampling constraint [9, 10, 11] to ensure similar selection probabilities for all clients. However, this uniform approach may compromise selection performance, as it disregards the heterogeneous data distribution among clients. In contrast, our proposed fairness constraint specifically targets clients with similar data distributions. Thus, it allows for varying selection frequencies among clients with heterogeneous data distributions, such as clients 1 and 4 in Fig. 1. This flexibility enables the individual constraint to accommodate the data heterogeneity without compromising selection performance in single rounds.
Limitation of Prior Works. To address the client selection problem, a common approach is the random selection strategy [1]. Recent works have proposed various improvements, including contribution-based [12, 1, 13, 14, 15], loss-based [16, 17, 18], and cluster-based methods [19, 20]. However, these methods typically evaluate each round individually and fail to select clients from a long-term perspective, potentially leading to model biases and fairness issues. Additionally, the clustering pattern in data distributions may be unknown or non-existent, making the application of cluster-based methods challenging. Furthermore, fair federated learning has discussed collaborative fairness [13, 15, 21], group fairness [22, 22] and performance fairness [23, 24], where performance fairness aligns most closely with our objectives. Existing fair federated learning approaches often utilize a uniform sampling constraint [9, 10, 11] to ensure similar selection probabilities for all clients. However, it disregards the heterogeneous data distribution among clients and requires their selection probabilities to be uniformly similar. This may hinder the identification of the optimal subset of clients that best approximates the full set of clients.
In this study, we propose a novel long-term client selection strategy in federated learning (LongFed), designed to emulate the model performance achieved with full client participation. In a single round, we perform client selection by minimizing the estimation error between a representative subset of clients and the full set of clients. In multi-round selection, we propose a novel individual fairness constraint to promote the diversity of selected clients without compromising selection performance in single rounds. Subsequently, we formulate the client selection problem as a time-averaged optimization function and simplify it using Lyapunov optimization from control theory. In addition, leveraging the submodular nature of the optimization function, we propose a fast greedy algorithm to find the optimal subset of clients in each round. The convergence of the proposed strategy is also analyzed. We conduct extensive experiments to validate the proposed LongFed strategy. First, the proposed strategy effectively guides the system to converge along a trajectory similar to that of full client participation. Second, it outperforms prior works in terms of both accuracy and fairness. Last, the proposed strategy introduces minimal time overhead compared to the random selection strategy. Our contributions are as follows.
  • We propose a novel long-term client selection strategy designed to guide the system to converge along a trajectory similar to that of full client participation. This strategy does not require extensive assumptions on the data distribution.
  • We identify and address an issue that has been rarely discussed in previous works, i.e., the frequency of selecting clients with similar distributions. We propose an individual fairness criterion to mathematically formulate this issue and employ Lyapunov optimization to solve it.
  • We provide an extensive theoretical analysis of the convergence ability of the proposed strategy.
II Related Work
Existing literature relevant to our work can be broadly categorized into two groups: client selection methods and fairness issues in federated learning.
II-A Client Selection in Federated Learning
In vanilla federated learning systems [1], the random selection strategy is commonly employed to choose clients. Recent works have proposed various improvements, including contribution-based [12, 1, 13, 14, 15], loss-based [16, 17, 18], and cluster-based methods [19, 20].
The first type focuses on evaluating client contributions, where clients with higher contributions are assigned higher selection probabilities. One commonly used metric is the local data size, and clients with larger datasets are considered to have higher contributions [12, 1]. Another approach involves employing the Shapley value [25] from game theory, which calculates the average marginal model improvement of each client over all possible coalitions [13, 14, 15]. Clients that result in larger model performance improvements are regarded as making larger contributions and are assigned higher probabilities of selection. Additionally, some methods evaluate the similarity between the local model at each client and the aggregated global model at the server. Clients with higher similarities are considered to bring little improvement to the global model and are thus assigned lower selection probabilities in subsequent training epochs [21, 26].
The second type involves selecting clients based on their training losses. These methods consider that clients with higher training losses may struggle to effectively fit their local data. As a result, they assign higher selection probabilities or weights to these clients [16, 17]. FedCor is a representative work in loss-based approaches that employs Gaussian processes to model the loss correlations between clients and selects clients with a substantial reduction in expected global loss [18]. Furthermore, recognizing that clients may contain similar and redundant information, a diverse strategy is proposed to choose a subset of clients that can best represent the full client set [27].
The third approach involves a cluster-based strategy. Recognizing that clients may have different data distributions or have their own learning tasks, these clustering-based approaches divide the clients into several clusters. Clients within the same cluster exhibit similar data distributions, whereas those in different clusters may display significant variations in their data distributions. In aggregation, the server randomly selects a client from each cluster. In this cluster-based paradigm, FedCG leverages a graph neural network to capture gradient sharing across multiple clusters [19]. Moreover, IFCA addresses the cluster identification problem by determining the cluster membership of each client and optimizes each of the cluster models in a federated learning framework [20].
However, most of these methods select clients by evaluating each round individually, thus failing to optimize a series of client subsets from a long-term perspective. These limitations could cause potential model biases and fairness issues, ultimately resulting in lower convergence efficiency. Furthermore, if the clustering pattern among clients is known, the long-term client selection problem becomes straightforward: simply selecting one client randomly from each cluster in every communication round [20]. However, in practical scenarios, the clustering pattern is usually unknown or, in some cases, does not exist. This limitation renders cluster-based methods ineffective for solving the long-term client selection problem.
II-B Fairness in Federated Learning
There are three types of fairness in federated learning: collaborative fairness, group fairness, and performance fairness. Collaborative fairness emphasizes that clients who make larger contributions should be rewarded with correspondingly larger rewards, and the assessment of client contributions is a key challenge. Client contribution evaluation methods have been discussed in Section II-A in the context of contribution-based client selection methods.
Group fairness [28], also known as algorithmic fairness, emphasizes that model outputs should not unfairly discriminate against vulnerable or underrepresented groups, such as minorities, women, or the aged [29]. FairFed [22] serves as one of the representative works of addressing this concern. In FairFed, both global fairness metrics at the server and local fairness metrics at each client are defined. These metrics assess disparities in opportunities among different groups. FairFed then dynamically adjusts aggregation weights at each round based on the discrepancies between global and local fairness metrics, which effectively mitigates the biases towards sensitive attributes [22].
Performance fairness refers that the model should produce similar performance across all clients, where the uniformity of performance is commonly used as the metric for evaluation [30, 31]. Various approaches have emerged to achieve performance fairness. For example, in [24], fairness is defined by assigning uniform weight to each client and introduced as an additional constraint in the optimization function. Besides, [23] identifies a trade-off between model robustness and fairness, and proposes a personalized framework to inherently achieve both fairness and robustness benefits. Moreover, several works study the fairness issue by considering that the probability of being selected is similar for all clients [9, 10, 11].
However, the three types of fairness described above cannot effectively address the fairness issue outlined in Principle II. While collaborative fairness and group fairness focus on different aspects compared to our work, performance fairness aligns most closely with our objectives. In these works [9, 10, 11], the uniform sampling method ensures a similar probability of being selected for all clients. However, we argue that it is unsuitable for achieving the long-term client selection strategy. This is because the uniform sampling method disregards the heterogeneous data distribution among clients. Consequently, it may hinder the identification of the optimal subset of clients that best approximates the data distribution of full client participation.
III The Proposed Optimization Function
In this section, we begin with an introduction to the federated learning system. Next, we illustrate the client selection in a single round and across multiple rounds, followed by the proposed long-term optimization function. Last, we apply Lyapunov optimization to transform and simplify the optimization function.
III-A Preliminary of Federated Learning
In our work, we consider that the federated learning system consists of a central server and a set of clients denoted as =$\mathbb{N}=\{1,\cdots,N\}$. Each client $i\in\mathbb{N}$ has its own local dataset $\mathcal{D}_{i}$ with a size of $|\mathcal{D}_{i}|$. Due to limited communication bandwidth, in the $t$-th round, the server typically selects a subset of clients, denoted as $\mathbb{S}^{t}$ with $|\mathbb{S}^{t}|=K<N$. Clients within the subset $\mathbb{S}^{t}$ receive the global model $\bm{w}^{t}$ from the server. They then compute local updates on their respective local datasets, transmitting the resulting local gradients $\nabla f_{j}(\bm{w}^{t})$ back to the server. The server aggregates these gradients and updates the model parameters using
$\bm{w}^{t+1}=\bm{w}^{t}-\eta_{t}\sum_{j\in\mathbb{S}^{t}}\theta^{t}_{j}\nabla f_{j}(\bm{w}^{t}).$ (1)
Here, $\eta_{t}$ represents the predefined learning rate, and $\theta^{t}_{j}$ denotes the weight assigned to client $j\in\mathbb{S}^{t}$ in the $t$-th round. The training process lasts for $T$ communication rounds until the global model achieves convergence. Here, the subset $\mathbb{S}^{t}$ and the weight $\theta^{t}_{j}$ with $j\in\mathbb{S}^{t}$ are the variables to be determined.
III-B The Optimization Function
III-B1 Client Selection in a Single Round
In a single round, we identify the subset of clients that best approximates the data distribution of the full set of clients. Specifically, we evaluate the estimation error between the aggregated gradient obtained from the client subset $\mathbb{S}^{t}$ and the aggregated gradient obtained from the full client set $\mathbb{N}$. A small estimation error indicates that the subset of selected clients can effectively represent the data distribution of the full client set. Mathematically, we formulate it as
$\text{D}(\mathbb{S}^{t})=\min_{\theta_{j}^{t}>0}||\sum_{i\in\mathbb{N}}\nabla f_{i}(\bm{w}^{t})-\sum_{j\in\mathbb{S}^{t}}\theta_{j}^{t}\nabla f_{j}(\bm{w}^{t})||_{2}^{2}.$ (2)
Here, $\nabla f_{i}(\bm{w}^{t})$ represents the gradient of the $i$-th client, $\sum_{i\in\mathbb{N}}\nabla f_{i}(\bm{w}^{t})$ is the aggregated gradient on the full client set $\mathbb{N}$, and $\sum_{j\in\mathbb{S}^{t}}\theta_{j}^{t}\nabla f_{j}(\bm{w}^{t})$ denotes the weighted sum of gradients on the client subset $\mathbb{S}^{t}$. Then, selecting a subset of clients that best approximates the data distribution of the full client set is equivalent to choosing a subset that minimizes the estimation error defined in Eq. (2).
Notice that the formation in Eq. (2) is similar to the concept of data coreset introduced in [32]. The data coreset involves selecting a weighted small sample of training data to approximate the gradient of the whole training data set. It is important to highlight that data coreset is commonly employed to enhance training efficiency in centralized machine learning scenarios [33, 34], while our focus is on optimizing client selection to reduce communication bandwidth in decentralized and federated scenarios. Moreover, data coreset is typically utilized for a one-time data selection process, while client selection in federated learning is a long-term process carried out across multiple rounds. The long-term optimization nature introduces new challenges, such as fairness issues and potential model biases introduced in Section I.
In Eq. (2), when clients are not selected in the $t$-th round, the gradients $\nabla f_{i}(\bm{w}^{t})$ for these clients become unknown, rendering the optimization of $\text{D}(\mathbb{S}^{t})$ intractable. To address this challenge, we obtain the upper bound of $\text{D}(\mathbb{S}^{t})$, denoted as $\text{DUB}(\mathbb{S}^{t})$ in Eq. (3), and minimize this upper bound instead. The proof of Theorem 1 is provided in the supplementary file.
Theorem 1.
Define
$\text{DUB}(\mathbb{S}^{t})\triangleq\sum_{i=1}^{N}\min_{j\in\mathbb{S}^{t}}\|\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})\|_{2}^{2},$ (3)
then $\text{DUB}(\mathbb{S}^{t})$ serves as an upper bound for $\text{D}(\mathbb{S}^{t})$.
Specifically, following the analysis in [32], we assume a mapping $\xi^{t}:\mathbb{N}\rightarrow\mathbb{S}^{t}$ that assigns each client $i\in\mathbb{N}$ to a client $j\in\mathbb{S}^{t}$, i.e., $\xi^{t}(i)=j$, indicating that client $i$ can be approximately represented by client $j$ in the $t$-th round. For a client $j\in\mathbb{S}^{t}$, let $\mathbb{C}^{t}_{j}=\{i\in\mathbb{N}\,|\,\xi^{t}(i)=j\}$ be the set of clients that can be represented by client $j$. The value $\theta^{t}_{j}=|\mathbb{C}^{t}_{j}|$ is the number of such clients being represented, and is used as the weight of client $j$ in Eq. (1). Based on the mapping $\xi^{t}$, we obtain the upper bound of the estimation error in Eq. (2), as stated in Theorem 1. In the later section, we utilize $\text{DUB}(\mathbb{S}^{t})$ when minimizing the estimation error $\text{D}(\mathbb{S}^{t})$ is required.
To manage the computation of $\text{DUB}(\mathbb{S}^{t})$ in Eq. (3), we introduce the metric $\text{Dist}_{i,j}(t)$ to quantify the distance between two clients $i$ and $j$ in the $t$-th round. This metric is defined as
$\text{Dist}_{i,j}(t)=||\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})||_{2}^{2}.$ (4)
Note that when clients $i,j\in\mathbb{N}$ are not selected in the $t$-th round, their gradients $\nabla f_{i}(\bm{w}^{t})$ and $\nabla f_{j}(\bm{w}^{t})$, and consequently $\text{Dist}_{i,j}(t)$, cannot be calculated. To address this issue, we propose partial updates for $\text{Dist}_{i,j}(t)$. Specifically, when $t=0$, all clients are selected, and we compute $\text{Dist}_{i,j}(t)$ for each pair of clients. Then, for $t\geq 1$, the server updates $\text{Dist}_{i,j}(t)$ using the gradients only from the selected $K$ clients. That is expressed as
$\text{Dist}_{i,j}(t)=\begin{cases}&\|\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})\|_{2}^{2},\;\text{if}\;i,j\in\mathbb{S}^{t},\\ &\text{Dist}_{i,j}(t-1),\;\text{otherwise}.\end{cases}$ (5)
Although using $\text{Dist}_{i,j}(t-1)$ to approximate $\text{Dist}_{i,j}(t)$ may introduce biases when either client is not selected, experimental results demonstrate that there are minimal impacts on the model convergence. This can be attributed to two reasons. First, there are limited variations in gradients between successive communication rounds. Moreover, the individual fairness constraint proposed in Section III-B2 ensures that each client has the opportunity of being selected and the corresponding $\text{Dist}_{i,j}(t)$ can be updated. In the following, we use $\text{Dist}_{i,j}(t)$ in Eq. (5) to replace the term $\|\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})\|_{2}^{2}$ in Eq. (3) to make $\text{DUB}(\mathbb{S}^{t})$ solvable.
III-B2 Client Selection Across Multiple Rounds
In our work, we propose a novel fairness constraint to guide the client selection process across multiple rounds. This constraint asserts that clients with similar data distributions should have similar frequencies of being selected.
To formulate the fairness constraint, we establish metrics to measure the similarity between a pair of clients in terms of their data distributions and the frequencies of being selected, respectively. First, we use $\text{Dist}_{i,j}(t)$ in Eq. (4) to measure the similarity of the data distribution and use Eq. (5) to compute its value. Second, let $x_{i,t}$ represent whether the $i$-th client is selected in the $t$-th round, with $x_{i,t}=1$ if $i\in\mathbb{S}^{t}$ and $x_{i,t}=0$ otherwise. We use
$p_{i}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}(x_{i,t}).$ (6)
to evaluate the frequency of a client $i$ being selected throughout the total $T$ rounds. Note that $\{x_{i,t}\}$ forms a stochastic process, and the expectation operation $\mathbb{E}$ is applied in Eq. (6). Given two clients $i,j\in\mathbb{N}$, we use $|p_{i}-p_{j}|$ to quantify the difference between their frequencies of being selected.
Building on $\text{Dist}_{i,j}(t)$ and $p_{i}$, we employ the $\epsilon\text{-}\delta$-IF framework utilized in [35, 36] to quantify the fairness constraint. We evaluate a client selection strategy at the end of the $T$-th communication round. Specifically, given $\epsilon,\delta\geq 0$, a client selection strategy is of individual fairness if for every pair of clients $i,j\in\mathbb{N}$ with $\text{Dist}_{i,j}(T)\leq\epsilon$, the difference in their frequencies of being selected, namely, $p_{i}$ and $p_{j}$, should not exceed $\delta$. The mathematical formulation of $\epsilon\text{-}\delta$-IF is provided in Definition 1, and the impact of parameter selection for $\epsilon$ and $\delta$ is discussed in Section V-C.
Definition 1.
($\epsilon\text{-}\delta$-IF). Consider $\epsilon,\delta\geq 0$, a client selection strategy is said to be of individual fairness if
$\forall\,i,j\in\mathbb{N},\,\,\text{Dist}_{i,j}(T)\leq\epsilon\,\,\Rightarrow\,\,|p_{i}-p_{j}|\leq\delta.$ (7)
Note that $\epsilon\text{-}\delta$-IF requires examining Eq. (7) for all pairs of clients, leading to significant computational overhead and rendering subsequent optimization challenging. To address this issue, we propose to determine a reference client $i^{\star}$ for each client $i$, which exhibits the largest difference in the frequency of being selected among clients whose gradient distances are less than $\epsilon$. That is,
$i^{\star}=\text{argmax}_{\text{Dist}_{i,j}(T)\leq\epsilon}\,|p_{i}-p_{j}|,\;\forall\;j\in\mathbb{N}.$ (8)
Then, given a client $i$ and its reference client $i^{\star}$, a client selection strategy is certified as $\epsilon\text{-}\delta$-IF if
$|p_{i}-p_{i^{\star}}|\leq\delta,\quad\forall\;i\in\mathbb{N}.$ (9)
By introducing the reference client in Eq. (8), the evaluation of $\epsilon\text{-}\delta$-IF is simplified from examining Eq. (7) for pairs of clients $i,j\in\mathbb{N}$ to evaluating Eq. (9) for individual clients $i\in\mathbb{N}$. This reduction in the number of variables facilitates further optimization.
III-B3 The Optimization Function
Building on the estimation error in Eq. (2) and the individual fairness constraint in Eq. (9), we formulate the client selection strategy as an optimization problem. The objective is to select a series of subsets $\{\mathbb{S}^{1},\cdots,\mathbb{S}^{T}\}$ with $|\mathbb{S}^{t}|=K$ that minimize the expected estimation error over all $T$ rounds while adhering to the individual fairness constraint. That is,
(P1) $\min_{\{\mathbb{S}^{1},\cdots,\mathbb{S}^{T}\}}\;\lim_{T\to+\infty}\frac{1}{T}\,\sum_{t=1}^{T}\mathbb{E}\left[\text{DUB}(\mathbb{S}^{t})\right],$ (10)
[…]
However, directly solving the optimization function in Eq. (10) is infeasible. The optimization objective and constraints, particularly $p_{i}$ in Eq. (9), are presented in a time-averaged form. That is, these values are determined by averaging over all $T$ rounds, which can only be accomplished at the end of training. In contrast, federated learning requires clients to be selected online in each round. This misalignment poses a practical implementation challenge. In our work, the solution to address this misalignment is proposed in Section III-C.
III-C Problem Transformation Under Lyapunov Optimization
To solve P1 in Eq. (10), we leverage Lyapunov optimization, a technique from control theory used to analyze the stability of dynamic systems [37, 9]. The main idea behind Lyapunov optimization is to break down the long-term time-averaged constraints in the optimization problem into constraints that can be adhered to in each communication round. This allows us to optimize the objective function directly in each round without violating the constraints specific to that round. Furthermore, by ensuring that these constraints are adhered to in each round, we can automatically achieve compliance with the long-term time-averaged constraints at the end of the $T$ round. By leveraging Lyapunov optimization, the problem stated as P1 in Eq. (10) is ultimately converted into the problem P3 in Eq. (23), which can be efficiently solved online in each communication round. Details of problem transformation are described below.
Before employing Lyapunov optimization, we remove the absolute value sign from the constraint in Eq. (9) and rephrase it as two equivalent constraints,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}(x_{i,t}-x_{i^{\star},t})-\delta\leq 0,\;\forall\,i\in\mathbb{N},$ (11)
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}(-x_{i,t}+x_{i^{\star},t})-\delta\leq 0,\;\forall\,i\in\mathbb{N}.$ (12)
Lyapunov optimization typically involves three steps [37]. First, a virtual queue is defined for each time-averaged constraint. This step converts the time-averaged constraints into stability constraints for virtual queues. Second, a Lyapunov function is defined based on virtual queues, and the Lyapunov drift is introduced. This step breaks down the time-averaged constraints into requirements for Lyapunov drift in each communication round. Finally, the Lyapunov drift is integrated into the optimization objective, allowing us to minimize both the Lyapunov drift and the optimization objective simultaneously. In the following, we illustrate how to transform problem P1 using these three steps.
First, we define two virtual queues, namely $Z_{i}(t)$ for the constraint in Eq. (11) and $Q_{i}(t)$ for the constraint in Eq. (12), for all clients $i\in\mathbb{N}$. Specifically, these queues are initialized as $Z_{i}(0)=0$ and $Q_{i}(0)=0$, and updated according to the following rule
$Z_{i}(t+1)=\max\left\{Z_{i}(t)+x_{i,t}-x_{i^{\star},t}-\delta,0\right\},$ (13)
$Q_{i}(t+1)=\max\left\{Q_{i}(t)-x_{i,t}+x_{i^{\star},t}-\delta,0\right\}.$
By the introduction of $Z_{i}(t)$ and $Q_{i}(t)$, we have Theorem 2, which converts the long-term constraints on $x_{i,t}$ in Eq. (11) and (12) into the stability constraints for $Z_{i}(t)$ and $Q_{i}(t)$. The proof of Theorem 2 is provided in the supplementary file.
Theorem 2.
The constraints in Eq. (11) and (12) hold if $Z_{i}(t)$ and $Q_{i}(t)$ remain stable, that is,
$\lim_{T\to+\infty}\frac{\mathbb{E}[Z_{i}(T)]}{T}=0,\;\text{and}\;\lim_{T\to+\infty}\frac{\mathbb{E}[Q_{i}(T)]}{T}=0.$ (14)
To make $Z_{i}(t)$ and $Q_{i}(t)$ remain stable, an intuitive approach is to bound the increase of these queues in each round so that they could not grow to infinity. Based on this intuition, in our work, we define a global queue $\Theta(t)$, which stores the state of $Z_{i}(t)$ and $Q_{i}(t)$ for all clients. That is,
$\Theta(t)\triangleq\left[Z_{1}(t),\cdots,Z_{n}(t),Q_{1}(t),\cdots,Q_{n}(t)\right].$ (15)
Second, following Lyapunov optimization, we define the Lyapunov function [37] as
$L\left(\Theta(t)\right)\triangleq\frac{1}{2}\sum_{i=1}^{N}\left[Z_{i}^{2}(t)+Q_{i}^{2}(t)\right],$ (16)
which represents the sum of the squares of all elements in $\Theta(t)$. Then, the increase of $\Theta(t)$ from the communication round $t$ to $(t+1)$ is formulated as
$\Delta\left(\Theta(t)\right)\triangleq\mathbb{E}\left[L\left(\Theta(t+1)\right)-L\left(\Theta(t)\right)|\Theta(t)\right],$ (17)
which is called Lyapunov drift [37]. If the drift $\Delta\left(\Theta(t)\right)$ remains sufficiently small in each round, the constraints in Eq. (11) and Eq. (12) will be satisfied after $T$ rounds. That is, introducing $\Delta\left(\Theta(t)\right)$ allows us to break down the time-averaged constraints in Eq. (11) and Eq. (12) into the specific requirements for $\Delta\left(\Theta(t)\right)$ in each communication round.
At last, to ensure that $\Delta\left(\Theta(t)\right)$ remains sufficiently small, a straightforward method to combine the optimization objection with the drift $\Delta\left(\Theta(t)\right)$, and minimize both simultaneously. Mathematically, in our work, the optimization problem in Eq. (10) is transformed into
[…]
with $|\mathbb{S}^{t}|=K$ as a constraint, where $V$ is a predefined trade-off factor. Note that the expectation notation $\mathbb{E}$ is dropped since we are only concerned with a single communication round.
However, solving P2 in Eq. (18) still presents two challenges. First, determining $i^{\star}$ requires the information about $p_{i}$, which is not available until the end of training. To address the issue, we calculate the frequency of a client $i$ being selected up to the $t$-th round, denoted as $p_{i}(t)=\frac{1}{t}\sum_{k=1}^{t}x_{i,t}$. Then, we determine the reference client in the $t$-th round by
$i^{\star}_{t}=\text{argmax}_{\text{Dist}_{i,j}(t)\leq\epsilon}|p_{i}(t)-p_{j}(t)|,\;\forall\,j\in\mathbb{N}.$ (19)
In our experiment, we replace the reference client $i^{\star}$ in Eq. (8) with $i^{\star}_{t}$ in Eq. (19), thus it can be determined using information available up to the $t$-th round.
Additionally, because the computation of $\Delta\left(\Theta(t)\right)$ requires information about $L\left(\Theta(t+1)\right)$, which is not available in the current round $t$. To overcome this issue, we derive an upper bound for $\Delta\left(\Theta(t)\right)$ using Theorem 3 and minimize the upper bound instead. The proof is provided in the supplementary file.
Theorem 3.
Define
$m_{i,t}=x_{i,t}-x_{i^{\star}_{t},t}-\delta,\;\text{and}\;n_{i,t}=-x_{i,t}+x_{i^{\star}_{t},t}-\delta,$ (20)
then $\Delta\left(\Theta(t)\right)$ is bounded by
$\Delta\left(\Theta(t)\right)\leq B+\sum_{i=1}^{N}\left[Z_{i}(t)m_{i,t}+Q_{i}(t)n_{i,t}\right],$ (21)
where $B$ is a constant.
Based on Theorem 3, the problem P2 in Eq. (18) is transformed into
$\min_{\mathbb{S}^{t}}\;(1-V)\sum_{i=1}^{N}\left[Z_{i}(t)m_{i,t}+Q_{i}(t)n_{i,t}\right]+V\cdot\text{DUB}(\mathbb{S}^{t}).$ (22)
Substituting $\text{DUB}(\mathbb{S}^{t})$ by Eq. (3) and moving the sum sign outside of the minimum sign, the problem in Eq. (22) is equivalent to
$\min_{\mathbb{S}^{t}}\;G(\mathbb{S}^{t})=\sum_{i=1}^{N}\min_{j\in\mathbb{S}^{t}}\Big{\{}(1-V)\big{[}Z_{i}(t)m_{i,t}+Q_{i}(t)n_{i,t}\big{]}$ (23)
[…]
where $Z_{i}(t)$ and $Q_{i}(t)$ are defined in Eq. (13), $m_{i,t}$ and $n_{i,t}$ are defined in Eq. (20), and $V$ is a hyperparameter that balances the optimization function and the fairness constraint. With an appropriate $V$, the solution in Eq. (23) can infinitely approximate the optimal solution in Eq. (18) [38]. In the following, we select clients by minimizing $G(\mathbb{S}^{t})$ in Eq. (23) in each round.
IV The Proposed LongFed
IV-A The Client Selection Strategy
The optimization problem in Eq. (23) is NP-hard as it involves calculating the value of $G(\mathbb{S}^{t})$ for $\frac{N!}{K!(N-K)!}$ subsets, where !!! denotes the factorial function [39]. To address this issue, we exploit the submodular nature of $G(\mathbb{S}^{t})$.
Specifically, a set function $g:2^{\mathbb{N}}\rightarrow\mathbb{R}$ is submodular if for every $A\subseteq B\subseteq\mathbb{N}$ and $\𝐵i\in\mathbb{N}\backslash B$ it holds that $g(A\cup\{i\})-g(A)>g(B\cup\{i\})-g(B)$. One typical example of a submodular function is the facility location function [40]. Suppose we aim to select locations from a set of positions =$\mathbb{N}=\{1,\ldots,N\}$ to open facilities and to serve a collection of $K$ users. If a facility is located at position $j$, the service it provides to user $i$ is quantified by $M_{i,j}$. Each user is assumed to select the facility with the highest service, and the total service provided to all users is modeled by the set function
$f(\mathbb{S})=\sum_{i=1}^{m}\max_{j\in\mathbb{S}}M_{i,j},$ (24)
where $f(\emptyset)=0$. If $M_{i,j}\geq 0$ for all $i,j$, then $f(\mathbb{S})$ is a monotone submodular function. By introducing an auxiliary element $e$, the set function $G(\mathbb{S}^{t})$ in Eq. (23) is transformed into
$\overline{G}(\mathbb{S}^{t})=G(\{e\})-G(\mathbb{S}^{t}\cup\{e\}),$ (25)
which is a facility location function and has a submodular nature. $\overline{G}(\mathbb{S}^{t})$ measures the decrease in the value of $G(\mathbb{S}^{t})$ associated with the set $\mathbb{S}^{t}$ compared to that associated with just the auxiliary element $e$. Without loss of generality, we set the auxiliary element as $e=1$. Consequently, minimizing $G(\mathbb{S}^{t})$ in Eq. (23) is equivalent to maximizing $\overline{G}(\mathbb{S}^{t})$ in Eq. (25).
Algorithm 1 The Proposed Client Selection Strategy
1:Input: $Z_{i}(t)$, $Q_{i}(t)$
2:Output: the selected subset $\mathbb{S}^{t}$
3:Initialize $\mathbb{S}_{0}^{t}=\emptyset$, $\mathbb{P}_{0}=\mathbb{N}$, and $e=1$.
4:for  $k\in[0,K-1]$ do
5:     Determine the reference client $i^{\star}_{t}$ for $i\in\mathbb{P}_{k}$
6:     Compute $m_{i,t}$ and $n_{i,t}$ for $i\in\mathbb{P}_{k}$
7:     Calculate $G(\mathbb{S}_{k}^{t}\cup\{i\}),\forall\,i\in\mathbb{P}_{k}$
8:     Identify the client $i_{\text{max}}=\text{argmax}_{i}\overline{G}(\mathbb{S}_{k}^{t}\cup\{i\})$
9:     $\mathbb{S}_{k+1}^{t}\leftarrow\mathbb{S}_{k}^{t}\cup i_{\text{max}},\;\mathbb{P}_{k+1}\leftarrow\mathbb{P}_{k}\backslash i_{\text{max}}$
10:end for
Prior studies show that the greedy algorithm is an effective solution for finding the maximum value of a submodular function [41, 39]. Following this greedy algorithm, we propose a strategy to select clients in the $t$-th round, as outlined in Algorithm 1. The proposed strategy starts with an empty set $\mathbb{S}_{0}^{t}=\emptyset$, and initializes a candidate set as $\mathbb{P}_{0}=\mathbb{N}$ (Line 3). In an iteration $k\in[0,K-1]$, we first determine the reference client $i^{\star}_{t}$ for client $i\in\mathbb{P}_{k}$, and compute $m_{i,t}$ and $n_{i,t}$ with $x_{i,t}=1$ if $i\in\mathbb{S}^{t}_{k}$, and similarly for $x_{i^{\star}_{t},t}$ (Line 5-6). Next, we calculate $\overline{G}(\mathbb{S}_{k}^{t}\cup\{i\})$ for all clients $i\in\mathbb{P}_{k}$, and identify the client $i_{\text{max}}$ with the maximum value (Line 7-8). Subsequently, the client $i_{\text{max}}$ is removed from $\mathbb{P}_{k}$ and added to the subset $\mathbb{S}_{k}^{t}$ (Line 9). This iteration continues until $K$ clients are selected. The complexity of the algorithm is $O(KN)$.
Based on the client selection strategy in Algorithm 1, we further present the proposed federated training algorithm, as outlined in Algorithm 2. First, the server initializes the queues $Z_{i}(0)$ and $Q_{i}(0)$, along with the model parameter $\bm{w}^{0}$ (Line 3). Subsequently, the server selects the subset of clients $\mathbb{S}^{t}$ (Line 5-9). If the communication round $t=0$, all clients are selected; otherwise, the server selects $K$ clients according to Algorithm 1. The server then sends the model parameter $\bm{w}^{t}$ to the selected clients, and these clients perform local training and send the gradients back to the server (Line 10-11). The server updates $Z_{i}(t)$ and $Q_{i}(t)$ according to Eq. (13), and aggregates these results to obtain the model parameter $\bm{w}^{t+1}$ according to Eq. (1). The iteration is repeated until completing the $T$ rounds.
Algorithm 2 The Federated Training Algorithm
1:Input: $\epsilon$, $\delta$, $V$ and $T$
2:Output: The trained model $\bm{w}^{T}$
3:Initialize $\bm{w}^{0}$, $Z_{i}(0)=Q_{i}(0)=0$,
4:for $t\in[0,T]$ do
5:     if $t=0$ then
6:         Select all clients with $\mathbb{S}^{t}=\mathbb{N}$
7:     else
8:         Select K clients $\mathbb{S}^{t}$ according to Algorithm 1
9:     end if
10:     The server sends $\bm{w}^{t}$ to the selected clients in $\mathbb{S}^{t}$
11:     Clients train local models in parallel and send the gradients $\nabla f_{i}(\bm{w}^{t})$ to the server
12:     The server update $Z_{i}(t+1)$, $Q_{i}(t+1)$
13:     The server aggregate the results and obtain $\bm{w}^{t+1}$
14:end for
IV-B Convergence Analysis
We provide a theoretical analysis to demonstrate the convergence ability of the proposed client selection strategy. The analysis uses FedAvg as the aggregation method, and it can be extended to other federated optimization methods as well. To implement the theoretical analysis, we establish six assumptions regarding the local models and data distribution heterogeneity among clients.
First, based on the effective greedy selection strategy in Algorithm 1, we assume that the estimation error between the client subset and the full client set (Eq. (2)) is small and can be quantified by a variable $\rho$, as stated in Assumption 1. This assumption has also been studied in the context of data coreset selection [32] and client selection methods [27]. Note that $\rho$ is used as a measure to characterize the quality of the estimation, and the analysis holds for any $\rho<\infty$.
Assumption 1. At each communication round $t$, we assume that the gradient aggregated from the selected subset of clients can provide a good approximation of the gradient aggregated from the full set, i.e.,
$\left\|\sum_{i\in\mathbb{N}}\nabla f_{i}(\bm{w}^{t})-\sum_{j\in\mathbb{S}^{t}}\theta_{j}^{t}\nabla f_{j}(\bm{w}^{t})\right\|\leq\rho.$ (26)
Next, we outline the assumptions regarding local models $f_{1},\cdots,f_{N}$ and their gradients $\nabla f_{1}(\bm{w}^{t}),\cdots,\nabla f_{N}(\bm{w}^{t})$, as stated in Assumption 2-5. These assumptions are standard and widely used in the federated optimization literature [42, 43, 18, 27].
Assumption 2. $f_{1},\cdots,f_{N}$ are all $L$-smooth. Formally, for all $\bm{v}$ and $\bm{w}$, we have
$f_{k}(\bm{v})\leq f_{k}(\bm{w})+(\bm{v}-\bm{w})^{T}\nabla f_{k}(\bm{w})+\frac{L}{2}\|\bm{v}-\bm{w}\|_{2}^{2}.$ (27)
Assumption 3. $f_{1},\cdots,f_{N}$ are all $\mu$-strongly convex. Formally, for all $\bm{v}$ and $\bm{w}$, it holds
$f_{k}(\bm{v})\geq f_{k}(\bm{w})+(\bm{v}-\bm{w})^{T}\nabla f_{k}(\bm{w})+\frac{\mu}{2}\|\bm{v}-\bm{w}\|_{2}^{2}.$ (28)
Assumption 4. The variance of gradients for $f_{i}$ is bounded for $i\in\mathbb{N}$. Formally, letting $\alpha_{i}^{t}$ be a data sample randomly chosen from the local dataset of the client $i$, we have
$\mathbb{E}\left[\|\nabla f_{i}(\bm{w}_{i}^{t},\alpha_{i}^{t})-\nabla f_{i}(\bm{w}_{i}^{t})\|\right]\leq B_{2}.$ (29)
Assumption 5. The expected squared norm of gradients is uniformly bounded, that is,
$\mathbb{E}\left[\|\nabla f_{i}(\bm{w}_{i}^{t},\xi_{i}^{t})\|\right]\leq B_{3}.$ (30)
Furthermore, we introduce the term $\Gamma$ in Assumption 6 to quantify the data heterogeneity among clients. If the data distribution among clients is independently and identically distributed, then $\Gamma$ approaches zero as the number of clients grows. Conversely, if the data distribution is heterogeneous, the magnitude of $\Gamma$ reflects the degree of heterogeneity.
Assumption 6. Let $f^{\star}$ and $f_{i}^{\star}$ be the minimum values of $f$ and $f_{i}$, respectively. We consider the degree of data heterogeneity to be bounded, that is,
$\Gamma=\|f^{\star}-\sum_{i=1}^{N}\theta_{i}f_{i}^{\star}\|\leq B_{4}.$ (31)
Based on these assumptions, the proposed client selection strategy is demonstrated to converge to the global optimal parameter $\bm{w}^{\star}$ at a rate of $\mathcal{O}(1/t)$ for heterogeneous data settings, as stated in Theorem 4. The proof is provided in the supplementary file. The convergence rate $\mathcal{O}(1/t)$ is the same as existing client selection methods [27, 18, 43].
Theorem 4.
Under Assumptions 1-6, we have
$\mathbb{E}\|\bm{w}^{\star}-\bm{w}^{t}\|_{2}^{2}\leq\mathcal{O}(1/t)+\mathcal{O}(\rho).$ (32)
The term $\rho$ encodes the estimation error in Eq. (2). When more clients are selected in a communication round, the subset of selected clients can better approximate the data distribution of the full client set, then the term $\rho$ is smaller. Particularly, $\rho$ becomes zero when all clients are selected in each round, i.e., $K=N$. We also analyze the impact of varying $K$ empirically and propose an adaptive client selection method [44]. However, to reduce communication overhead, the number of selected clients is usually less than the total number of available clients, i.e., $K<N$. In this case, $\rho$ remains a non-vanishing term. In experiments, we observe that the proposed client selection strategy guides the system to converge along a trajectory similar to that of full client participation, but there exists a non-diminishing solution bias dependent on $\rho$. This observation is consistent with our theoretical analysis.
V Experiment
V-A Experimental Settings
We consider the cross-device federated learning scenario where a large number of clients exist, each with limited computational power. In our experiments, we fix the number of clients at $N=100$ and employ FedAvg [1] as the aggregation method. The proposed client selection strategy is evaluated on two datasets, FMNIST [45] and CIFAR-10 [46]. Following the approach in [18], we opt for basic and small-scale models to accommodate the clients’ restricted computational resources. Specifically, for the FMNIST dataset, we utilize a multilayer perception (MLP) with two hidden layers as the global model. For CIFAR-10, our global model adopts a convolutional neural network (CNN) architecture consisting of three convolutional layers, with the outputs of these layers fed into an MLP layer. Further details on training and hyperparameter settings can be found in the supplementary file.
We explore four data partitioning methods, including an independently and identically distributed (IID) approach and three heterogeneous approaches. In the IID approach, we randomly partition the dataset into $N$ parts, with each client assigned one part. We select $K=5$ clients in each round. Three heterogeneous data partitioning approaches are as follows.
(i) 1 Shard Per Client (1SPC). Following the setting in [1], we divide the dataset into $N$ shards, ensuring that data within a shard shares the same label. We randomly assign a shard to each client so that a client has data with only one label. This is the most heterogeneous case. In this case, we select $K=10$ clients in each round.
(ii) 2 Shards Per Client (2SPC). Similar to the 1SPC setting, the dataset is partitioned into $2N$ shards, where data within each shard shares the same label. Clients are randomly assigned two shards, where most clients have data with two labels (two shards with different labels), while a few clients have data with only one label (two shards with the same label). In this case, we set $K=5$ in each round.
(iii) Dirichlet Distribution (Dir). We partition the dataset based on a Dirichlet distribution parameterized by a concentration variable $\alpha$, where a smaller value of $\alpha$ indicates higher heterogeneity. In our work, we set $\alpha=0.8$. Under this setting, clients have data with all labels, but the data size of each label varies among clients. We determine the data size of each label for each client following the analysis in [47, 18]. In this case, we choose $K=5$ clients in each round.
Figure 2: Test accuracy on FMNIST and CIFAR-10 under the IID scenario.
Note that the 1SPC and 2SPC scenarios address label shift heterogeneity, while the Dir scenario accounts for heterogeneity in both labels and dataset size. Furthermore, apart from the assumption of a fixed number of clients ($K$) to be selected, we also explore the impact of varying $K$ and propose an adaptive method to determine its value. Details of the adaptive method can be found in [44].
Figure 3: Test accuracy on FMNIST and CIFAR-10 under three heterogeneous data partitioning settings.
TABLE I: Time overhead analysis on FMNIST in the 1SPC scenario.
Method Random AFL Powerd DFairFed
Time (ms) For Client Selection Only 0.038 0.153 0.016 0.529
Time (ms) For Completing a Communication Round 2019 2044 2033 2472
We conduct the experiments five times using different randomized initial conditions and present the average results. In our experiments, we set $\delta=0.1$ , $\epsilon=0.3$, and $V=0.9$ for FMNIST, and $\delta=0.05$, $\epsilon=0.05$, $V=0.7$ for CIFAR-10. We compare the proposed LongFed with three baseline methods: (1) the random selection strategy (Random) [1]; (2) active client selection strategy (AFL), where clients with higher training losses have higher probabilities of being selected [16]; (3) power-of-choice strategy (Powerd), where clients are selected based on the power of $d$ choices load balancing strategy [17]. These methods are commonly used as baselines in [18, 27]. Additionally, we compare our method with the full client participation method (Full) with $K=100$.
V-B Experimental Results
V-B1 Convergence Analysis
The experimental results under the IID scenario are presented in Fig. 2, and the results under three heterogeneous data partitioning methods are presented in Fig. 3. In Fig. 2 and Fig. 3, the $x$ axis represents the communication round, and the $y$ axis displays the test accuracy of the global model. For a clearer comparison, we apply the moving average rule [48] with a sliding window size of 10 to smooth the results.
Fig. 2 reveals that in the IID scenario, these methods achieve similar performance. For three heterogeneous scenarios on the FMNIST dataset, as shown in Fig. 3 (a), (b), and (c), respectively, the proposed method achieves performance that is the most similar to that with full client participation. Moreover, it exhibits faster convergence speed and superior test accuracy compared to prior works. Notably, there is an approximate $20\%$ improvement in the 1SPC scenario and an $8\%$ improvement in the 2SPC and Dir scenarios. Similar trends are also observed for the CIFAR-10 dataset.
Figure 4: Visualization of the selected clients on the FMNIST dataset under the 1SPC scenario. (a) is the t-SNE plot of client embeddings, where clients are organized into 10 clusters. (b)-(f) display the results of the selected clients using different client selection strategies, with the chosen clients marked by black stars.
Figure 5: Visualization of the selected clients on the FMNIST dataset under the 2SPC scenario. (a) shows the t-SNE plot of client embeddings, where the clustering pattern is not evident. In (b)-(f), circles are drawn around the selected clients, indicating that clients covered by a circle can be represented by the corresponding selected client.
V-B2 Time Overhead Analysis
We evaluate the time overhead of the proposed client selection strategy, and the results are presented in Table I. The time overhead analysis is conducted on the FMNIST dataset under the 1SPC scenario, with similar observations across other scenarios. In Table I, the first row focuses on the time solely for client selection within a communication round, while the second row denotes the overall time required for a complete communication round, including client selection, local updates, and global aggregation.
First, examining the time solely for client selection, our proposed method exhibits only a marginal increase (less than 0.4ms) compared to existing methods. This increment is acceptable in most real applications. In addition, the time required for client selection (about 0.5ms) is almost negligible compared to the total time required to complete a communication round (about 2000ms). This highlights that the majority of time is allocated to local updates within clients. Second, examining the time required to complete a round, our proposed strategy imposes only a modest additional computational time (about 400 ms) compared to existing approaches. This is primarily because evaluating the distance $\text{Dist}_{i,j}(t)$ in Eq. (5) is time-consuming, especially when the dimension of $\nabla f_{i}(\bm{w}^{t})$ is high. The time overhead can be reduced by employing a more efficient method for calculating distances in high-dimensional gradient space. Nevertheless, the marginal increase in time for the proposed strategy is justified when considering its superior performance in client selection.
V-B3 Visualization of Client Selection Strategy
We provide visualizations of the selected clients to offer an interpretable analysis of our proposed strategy. As shown in Fig. 4, in the 1SPC scenario on the FMNIST dataset, baseline methods typically select two or more clients from the same cluster. In contrast, the proposed LongFed, as shown in 4 (e), selects one client from each cluster, which effectively approximates the data distribution of the full client set. Additionally, as shown in Fig. 4 (f), it chooses a diverse subset of clients in the multi-round selection, adhering to Principle II.
The results in the 2SPC scenario are illustrated in Fig. 5, where the clustering pattern is not evident. For baseline methods, the circles around selected clients often overlap and cover only a minority of clients. However, in our proposed method, as shown in Fig. 5 (e), these circles cover the majority of clients. This suggests that the selected clients provide a better approximation of the full client set compared to the baseline approaches. Moreover, as shown in Fig. 5 (f), the strategy selects a diverse subset of clients across multiple rounds, satisfying Principle II. Visualization results of other scenarios are provided in the supplementary file.
V-C Individual Fairness Analysis
We analyze the proposed individual fairness in terms of the trade-off factor $V$, the similarity measure $\epsilon$, and the probability difference measure $\delta$. The results are illustrated using the 2SPC scenario on the CIFAR-10 dataset, and similar trends are observed for other scenarios. The random selection strategy is chosen as the baseline because it selects clients randomly and without bias. The impact of these parameters is assessed through two metrics: test accuracy and the standard deviation of the selection probability $p_{i}$ between clients with similar data distributions, denoted as $\sigma(t)$. To calculate $\sigma(t)$, we define $c_{i}(t)$ as the accumulated number of selections for client $i$ during previous $t$ communication rounds. That is, $c_{i}(t)=\sum_{k=1}^{t}x_{i,k}$. For each client $i$, we identify clients whose similarity with it is less than $\epsilon$, denoted as $\mathbb{I}_{i}(t)=\{j\in\mathbb{N}\mid\text{Dist}_{i,j}(t)<\epsilon\}$. We then compute the average selection count among clients in $\mathbb{I}_{i}(t)$ by
$\overline{c}_{i}(t)=\frac{1}{|\mathbb{I}_{i}(t)|}\sum_{j\in\mathbb{I}_{i}(t)}c_{j}(t),$ (33)
and the standard deviation in their selection counts by
$\sigma(t)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left[c_{i}(t)-\overline{c}_{i}(t)\right]^{2}}.$ (34)
A smaller $\sigma(t)$ indicates that the client selection strategy aligns more closely with the individual fairness constraint.
Trade-Off Factor $V$. We vary the value of $V$ from $0.0$, $0.4$, $0.6$, and $1.0$, and present the results in Fig. 6. First, we observe that the proposed method with these four values converges faster than the random strategy and achieves better performance. Among them, $V=0.6$ yields the best test accuracy. This suggests that the individual fairness constraint improves the overall performance. Second, the random strategy achieves the lowest standard deviation $\sigma(t)$. The proposed method exhibits significantly larger $\sigma(t)$ when $V=1.0$. This is because when $V=1.0$, our proposed method prioritizes minimizing the estimation error in Eq. (2) and neglects the fairness constraint. Furthermore, the standard deviation $\sigma(t)$ decreases significantly when $V\leq 0.6$. It indicates that the constraint in Eq. (7) indeed enhances the system fairness.
Similarity Measure $\epsilon$. We vary the value of $\epsilon$ across $0.05$, $0.1$, $0.2$, and $0.3$, and observe the corresponding test accuracy and standard deviation, as presented in Fig. 7. First, the results on test accuracy indicate that the proposed method outperforms the random strategy consistently across these $\epsilon$ values. On the other hand, a larger $\epsilon$ in the proposed method corresponds to a smaller standard deviation. Specifically, the results are comparable to the random selection strategy when $\epsilon>0.05$, and even surpass it when $\epsilon=0.3$. This is attributed to the fact that a larger $\epsilon$ enforces more clients to follow the fairness constraint, thus the standard deviation obtained through the proposed strategy gets closer to the random selection strategy.
Figure 6: Impact of the trade-off factor $V$.
Figure 7: Impact of the similarity measure $\epsilon$.
Probability Difference Measure $\delta$. We vary the value of $\delta$ across $0.01$, $0.1$, $0.15$, and $0.2$, and observe the corresponding test accuracy and standard deviation, as presented in Fig. 8. The results indicate that the proposed method consistently performs better than the random selection strategy across these $\delta$ values in terms of test accuracy. Moreover, a smaller $\delta$ in the proposed method corresponds to a smaller standard deviation. This is because a smaller $\delta$ narrows the range of probability difference between clients with similar distributions, resulting in a smaller standard deviation.
VI Conclusion
In our work, we introduce a novel long-term client selection approach for federated learning, aiming to emulate the optimization trajectory achieved by full client participation. We establish two guiding principles to steer the client selection process effectively. We formulate the client selection problem as a long-term optimization task, where the objective is to minimize the estimation error in gradient space between the selected subset of clients and the full client set. Besides, we impose the individual fairness constraint to ensure similar selection frequencies among clients with similar data distributions. Leveraging Lyapunov optimization techniques from control theory, we transform this problem and propose a greedy algorithm to find solutions efficiently. Experimental results demonstrate that our proposed strategy closely approximates the performance achieved under full client participation, with significant improvements over previous methods in terms of both accuracy and fairness. Additionally, our method exhibits efficiency by incurring minimal time overhead, which is often negligible when compared to local training duration.
Figure 8: Impact of the probability difference measure $\delta$.

### v2

@@
-Emulating Full Client Participation: A Long-Term Client Selection Strategy for Federated Learning
+Emulating Full Participation: An Effective and Fair Client Selection Strategy for Federated Learning
 Abstract
-Client selection significantly affects the system convergence efficiency and is a crucial problem in federated learning.
-Existing methods often select clients by evaluating each round individually and overlook the necessity for long-term optimization, resulting in suboptimal performance and potential fairness issues.
-In this study, we propose a novel client selection strategy designed to emulate the performance achieved with full client participation.
-In a single round, we select clients by minimizing the gradient-space estimation error between the client subset and the full client set.
-In multi-round selection, we introduce a novel individual fairness constraint, which ensures that clients with similar data distributions have similar frequencies of being selected.
-This constraint guides the client selection process from a long-term perspective.
-We employ Lyapunov optimization and submodular functions to efficiently identify the optimal subset of clients, and provide a theoretical analysis of the convergence ability.
-Experiments demonstrate that the proposed strategy significantly improves both accuracy and fairness compared to previous methods while also exhibiting efficiency by incurring minimal time overhead.
+In federated learning, client selection is a critical problem that significantly impacts both model performance and fairness.
+Prior studies typically treat these two objectives separately, or balance them using simple weighting schemes.
+However, we observe that commonly used metrics for model performance and fairness often conflict with each other, and a straightforward weighted combination is insufficient to capture their complex interactions.
+To address this, we first propose two guiding principles that directly tackle the inherent conflict between the two metrics while reinforcing each other.
+Based on these principles, we formulate the client selection problem as a long-term optimization task, leveraging the Lyapunov function and the submodular nature of the problem to solve it effectively.
+Experiments show that the proposed method improves both model performance and fairness, guiding the system to converge comparably to full client participation.
+This improvement can be attributed to the fact that both model performance and fairness benefit from the diversity of the selected clients’ data distributions.
+Our approach adaptively enhances this diversity by selecting clients based on their data distributions, thereby improving both model performance and fairness.
 Index Terms:
-Federated Learning, Client Selection, Individual Fairness, Coreset Selection, Lyapunov Optimization
+Federated Learning, Client Selection, Coreset Selection, Individual Fairness, Lyapunov Function
 I Introduction
-Federated learning (FL) facilitates collaborative model training without the necessity of sharing local data [1, 2], and is widely used in various domains [3, 4, 5].
-In FL, model parameters or gradient updates are frequently exchanged between the central server and participating clients, leading to substantial communication overhead.
-To address the challenge posed by the limited bandwidth, a common approach is to select a subset of clients for local training, which reduces the communication costs compared with involving all clients [6, 7].
-However, in heterogeneous scenarios, different selection strategies lead to distinct optimization trajectories, and an improper client selection strategy can cause the optimization trajectory to deviate seriously from the optimal one, significantly hindering the model performance.
-Therefore, a critical challenge is to develop an effective client selection strategy that reduces the communication overhead while ensuring model convergence.
-In most cases, the global models generated by full client participation are considered the gold standard or benchmark.
-This is because they represent models trained on the entire dataset available across all clients, offering a comprehensive reference point for evaluating model performance.
-Building upon this understanding, we propose an intuitive strategy where clients are selected in a manner that ensures the optimization trajectory closely resembles that achieved with full client participation over the long term.
-We refer to this strategy as the long-term client selection strategy.
-This strategy yields two key advantages.
-First, by emulating the optimization trajectory of full client participation, it aims to achieve comparable performance.
-Moreover, it reduces communication overhead by involving only a subset of clients in the process.
-Our Goal.
-In this work, we investigate the three questions: 1) what principles should be followed to implement this strategy; 2) how to mathematically formulate this strategy and provide a fast solution; and 3) is this strategy more effective than existing methods?
-To answer the first question, we propose two crucial principles to be followed.
-Principle I: In each round, the data distribution generated by the selected subset of clients should closely resemble the data distribution of the full client participation.
-For example, in Fig.
-1 (a), clients collaboratively train an image classification model, where each client has images of only one kind of animal, and three clusters exist in their data distributions.
-Due to limited bandwidth, the server is required only to select three clients in each round.
-In this case, as shown in Fig.
-1 (b), the ideal strategy is to select one client from each cluster instead of three clients from the same cluster.
-This ensures that the data distribution resulting from the subset of selected clients serves as a reliable estimation of that from the full set of clients.
-However, in real-world applications, the clustering pattern in data distributions may be unknown or non-existent.
-Therefore, the key challenge here is to determine the subset of clients that best approximates the full client set without imposing extensive assumptions on their data distributions.
-Figure 1: An example of the long-term client selection strategy.
-However, if only Principle I is followed and client selection is performed by evaluating each round individually, there is a risk that the same subset of clients may be consistently selected across multiple rounds, while others may have minimal opportunities for selection.
-This phenomenon is indeed observed in our experimental results.
-In such cases, the system may accurately predict outcomes for frequently selected clients but makes inaccurate or biased predictions for those rarely selected, resulting in degradation of the global model’s performance and ineffective convergence.
-Furthermore, these seldom-selected clients may perceive unfair treatment and lose motivation to participate, ultimately endangering the sustainability of the federated system.
-To mitigate this problem, we propose Principle II, which guides client selection across multiple rounds from a long-term perspective.
-Principle II: In multi-round selection, clients with similar data distributions should have similar frequencies of being selected.
-For example, as illustrated in Fig.
-1, clients 1, 2, and 3 should share similar selection frequencies, as should clients 4 and 5, while the frequencies for clients 1 and 4 may differ.
-This principle aligns with the concept of individual fairness (IF) in central machine learning [8], which posits that similar instances or individuals in a dataset should yield similar predictions or outcomes.
-In our work, we refer to this principle as the individual fairness constraint.
-Here, the key challenge is to quantify the similarity between clients in terms of data distribution and selection frequency, respectively, and further formulate individual fairness mathematically.
-This proposed individual fairness offers two key benefits.
-First, as illustrated in Fig.
-1 (c), it promotes the diversity of selected clients across multiple rounds and reduces model prediction bias.
-Furthermore, existing fair federated learning approaches often utilize a uniform sampling constraint [9, 10, 11] to ensure similar selection probabilities for all clients.
-However, this uniform approach may compromise selection performance, as it disregards the heterogeneous data distribution among clients.
-In contrast, our proposed fairness constraint specifically targets clients with similar data distributions.
-Thus, it allows for varying selection frequencies among clients with heterogeneous data distributions, such as clients 1 and 4 in Fig.
+Federated learning (FL) facilitates collaborative model training without the necessity of sharing local data [1, 2] and is widely used in various domains [3, 4].
+In FL, model parameters or gradient updates are frequently exchanged between the server and clients, which leads to substantial communication overhead.
+To address the challenge of limited bandwidth, a common approach is to select a subset of clients for local training [5, 6].
+Therefore, a critical challenge is how to select proper and representative clients to participate, which is known as the client selection problem.
+In heterogeneous scenarios—where clients hold training data with diverse distributions—the client selection problem becomes particularly critical.
+Specifically, client selection impacts the FL system in two key aspects.
+First, client selection influences model performance.
+Different selection strategies lead to distinct optimization trajectories, and an improper client selection may cause the optimization process to deviate significantly from the optimal path.
+Second, client selection affects fairness.
+The global model tends to perform better for frequently selected clients, as their data is better optimized, while producing biased predictions for less frequently selected clients.
+Since every client participates in the system with the expectation of obtaining accurate predictions for its own data, this selection imbalance may drive underrepresented clients to leave the system.
+Therefore, it is essential to establish an effective and fair client selection strategy.
+Limitation of Prior Works.
+Prior studies usually treat model performance and fairness as separate objectives and attempt to balance the two using simple weighting schemes [7, 8, 9].
+However, model performance and fairness metrics commonly in use are often conflicting, and a simple weighted combination is insufficient to capture their intricate interactions.
+Specifically, to achieve high model performance, existing client selection methods [10, 11, 12] often prioritize clients with higher training losses, as these clients are considered more challenging to fit their local data.
+However, such loss-guided methods can lead to biased predictions for clients that are rarely selected.
+On the other hand, to enhance fairness, existing approaches typically focus on performance fairness [13, 14, 7, 8, 9], which requires the model to deliver similar performance or uniform selection probabilities across all clients.
+However, this uniform selection approach overlooks the heterogeneous nature of clients’ data distributions, thereby sacrificing model performance.
+Therefore, a critical challenge lies in reconciling the conflicting goals of model performance and fairness, particularly in heterogeneous scenarios.
+Figure 1: Visualization of the client selection results.
+Clients are represented by blue dots, and selected clients are marked with orange stars.
+In the non-clustering scenario, clients arranged in a circle are effectively represented by the selected client positioned at the center.
+To design an effective and fair client selection strategy, we propose two guiding principles.
+These principles address the inherent conflict between the two metrics by capturing their interactions.
+First, from the perspective of model performance, we propose Principle I: The data distribution generated by the selected subset of clients should closely resemble the data distribution of the full client participation.
+Full client participation is chosen as the standard because models trained with all clients generally yield robust results, serving as a comprehensive benchmark—except in certain extreme cases where some clients possess highly noisy or corrupted data that can skew the model’s performance.
+Importantly, for clients that are less frequently selected in the loss-guided methods [10, 11, 12], this strategy ensures they are not excluded from selection by approximating the distribution of full client participation.
+Furthermore, from the perspective of fairness, we introduce Principle II: clients with similar data distributions should have similar frequencies of being selected.
+This principle aligns with the concept of Individual Fairness (IF) [15], which posits that instances with similar features in a dataset should yield similar predictions or outcomes.
+We refer to this principle as the individual fairness constraint.
+Unlike the uniform sampling strategies used in prior works [8, 9, 13], this constraint specifically focuses on clients with similar data distributions.
+It is better suited to heterogeneous data scenarios, as it avoids sacrificing model performance by preventing the uniform selection of clients who contribute minimally to model improvement.
+The two principles are not independent and mutually reinforce each other.
+Consider an extreme case where there are two clusters of clients, and the data distributions are similar within each cluster.
+If only Principle II is applied, it may result in the system selecting clients exclusively from one cluster while neglecting the other.
+This phenomenon was observed in our experimental results, whereas Principle I helps mitigate this issue.
+On the other hand, if only Principle I is applied, the system may end up selecting the same client from each cluster repeatedly, whereas Principle II prevents this bias.
+In this study, we integrate the two principles and propose an effective and fair client selection strategy for federated learning, called LongFed.
+We begin by providing a mathematical formulation for both Principle I and Principle II.
+Then, we model the client selection problem as a long-term optimization function, introducing a tradeoff factor to balance the two principles.
+To solve this optimization, we simplify it using the Lyapunov optimization from control theory, and propose a fast greedy algorithm based on the submodular nature of the problem.
+We also theoretically analyze the convergence of the proposed strategy, demonstrating that it converges at a rate of $\mathcal{O}(1/t)$, which is the same as loss-guided selection methods [10, 11, 12, 16].
+We evaluate the proposed strategy through extensive experiments, with results demonstrating that our method enhances both model performance and fairness.
+Regarding model performance, the strategy effectively guides the system to converge along a trajectory similar to that of full client participation, outperforming prior methods in most cases.
+Additionally, the strategy achieves strong fairness, evidenced by a low standard deviation in the selection frequencies of clients with similar data distributions.
+This improvement can be attributed to the fact that both model performance and fairness benefit from the diversity of the selected clients’ data distributions, which is promoted by our two proposed principles.
+Unlike existing fair federated learning approaches [8, 9, 13] that treat all clients equally, our method adaptively enhances this diversity based on clients’ data distributions, making it better suited for heterogeneous environments.
+Additionally, our method introduces only a marginal time increase (less than 0.4 ms) compared to existing approaches.
+We also provide visualization results to illustrate the effectiveness of our method.
+An example is shown in Fig.
 1.
-This flexibility enables the individual constraint to accommodate the data heterogeneity without compromising selection performance in single rounds.
-Limitation of Prior Works.
-To address the client selection problem, a common approach is the random selection strategy [1].
-Recent works have proposed various improvements, including contribution-based [12, 1, 13, 14, 15], loss-based [16, 17, 18], and cluster-based methods [19, 20].
-However, these methods typically evaluate each round individually and fail to select clients from a long-term perspective, potentially leading to model biases and fairness issues.
-Additionally, the clustering pattern in data distributions may be unknown or non-existent, making the application of cluster-based methods challenging.
-Furthermore, fair federated learning has discussed collaborative fairness [13, 15, 21], group fairness [22, 22] and performance fairness [23, 24], where performance fairness aligns most closely with our objectives.
-Existing fair federated learning approaches often utilize a uniform sampling constraint [9, 10, 11] to ensure similar selection probabilities for all clients.
-However, it disregards the heterogeneous data distribution among clients and requires their selection probabilities to be uniformly similar.
-This may hinder the identification of the optimal subset of clients that best approximates the full set of clients.
-In this study, we propose a novel long-term client selection strategy in federated learning (LongFed), designed to emulate the model performance achieved with full client participation.
-In a single round, we perform client selection by minimizing the estimation error between a representative subset of clients and the full set of clients.
-In multi-round selection, we propose a novel individual fairness constraint to promote the diversity of selected clients without compromising selection performance in single rounds.
-Subsequently, we formulate the client selection problem as a time-averaged optimization function and simplify it using Lyapunov optimization from control theory.
-In addition, leveraging the submodular nature of the optimization function, we propose a fast greedy algorithm to find the optimal subset of clients in each round.
-The convergence of the proposed strategy is also analyzed.
-We conduct extensive experiments to validate the proposed LongFed strategy.
-First, the proposed strategy effectively guides the system to converge along a trajectory similar to that of full client participation.
-Second, it outperforms prior works in terms of both accuracy and fairness.
-Last, the proposed strategy introduces minimal time overhead compared to the random selection strategy.
+When clients exhibit clear clustering patterns, the proposed strategy selects one client from each cluster and chooses different clients across multiple rounds.
+This selection process aligns with the clustered federated learning [17, 18].
+More importantly, in scenarios where the clustering pattern is not evident—which is more common in practical cases—the proposed strategy selects clients that cover a majority of client population.
+This diversity not only enhances model performance but also improves fairness.
 Our contributions are as follows.
-  • We propose a novel long-term client selection strategy designed to guide the system to converge along a trajectory similar to that of full client participation.
-This strategy does not require extensive assumptions on the data distribution.
-  • We identify and address an issue that has been rarely discussed in previous works, i.e., the frequency of selecting clients with similar distributions.
-We propose an individual fairness criterion to mathematically formulate this issue and employ Lyapunov optimization to solve it.
+  • We identify the inherent conflict between model performance and fairness in the client selection problem and propose a strategy that leads to improvements in both.
+  • We address a previously underexplored issue: the frequency of selecting clients with similar data distributions.
+We introduce an individual fairness criterion to mathematically formulate and effectively resolve this issue.
   • We provide an extensive theoretical analysis of the convergence ability of the proposed strategy.
@@
 Moreover, IFCA addresses the cluster identification problem by determining the cluster membership of each client and optimizes each of the cluster models in a federated learning framework [18].
-However, most of these methods select clients by evaluating each round individually, thus failing to optimize a series of client subsets from a long-term perspective.
-These limitations could cause potential model biases and fairness issues, ultimately resulting in lower convergence efficiency.
-Furthermore, if the clustering pattern among clients is known, the long-term client selection problem becomes straightforward: simply selecting one client randomly from each cluster in every communication round [20].
-However, in practical scenarios, the clustering pattern is usually unknown or, in some cases, does not exist.
-This limitation renders cluster-based methods ineffective for solving the long-term client selection problem.
 II-B Fairness in Federated Learning
@@
 FairFed then dynamically adjusts aggregation weights at each round based on the discrepancies between global and local fairness metrics, which effectively mitigates the biases towards sensitive attributes [28].
-Performance fairness refers that the model should produce similar performance across all clients, where the uniformity of performance is commonly used as the metric for evaluation [30, 31].
+Performance fairness refers that the model should produce similar performance across all clients, aligning most closely with our objectives.
 Various approaches have emerged to achieve performance fairness.
@@
 Moreover, several works study the fairness issue by considering that the probability of being selected is similar for all clients [8, 9, 29].
-However, the three types of fairness described above cannot effectively address the fairness issue outlined in Principle II.
-While collaborative fairness and group fairness focus on different aspects compared to our work, performance fairness aligns most closely with our objectives.
-In these works [9, 10, 11], the uniform sampling method ensures a similar probability of being selected for all clients.
-However, we argue that it is unsuitable for achieving the long-term client selection strategy.
-This is because the uniform sampling method disregards the heterogeneous data distribution among clients.
-Consequently, it may hinder the identification of the optimal subset of clients that best approximates the data distribution of full client participation.
+However, as introduced in Section I, the uniform selection method disregards the heterogeneous data distribution among clients.
+In summary, existing client selection methods only consider the model performance and do not address the fairness.
+Although some fair federated learning methods have been proposed, the uniform selection constraint adopted in their methods could disregard the heterogeneous data distribution among clients.
+Therefore, it remains a critical problem for client selection to simultaneously address model performance and fairness.
 III The Proposed Optimization Function
 In this section, we begin with an introduction to the federated learning system.
-Next, we illustrate the client selection in a single round and across multiple rounds, followed by the proposed long-term optimization function.
-Last, we apply Lyapunov optimization to transform and simplify the optimization function.
+Next, we provide a mathematical formulation for both Principle I and Principle II, and model the client selection problem as a long-term optimization function.
+Last, we apply Lyapunov optimization to simplify the optimization function.
 III-A Preliminary of Federated Learning
@@
 Each client $i\in\mathbb{N}$ has its own local dataset $\mathcal{D}_{i}$ with a size of $|\mathcal{D}_{i}|$.
-Due to limited communication bandwidth, in the $t$-th round, the server typically selects a subset of clients, denoted as $\mathbb{S}^{t}$ with $|\mathbb{S}^{t}|=K<N$.
+In the $t$-th round, the server selects a subset of clients, denoted as $\mathbb{S}^{t}$ with $|\mathbb{S}^{t}|=K<N$.
 Clients within the subset $\mathbb{S}^{t}$ receive the global model $\bm{w}^{t}$ from the server.
-They then compute local updates on their respective local datasets, transmitting the resulting local gradients $\nabla f_{j}(\bm{w}^{t})$ back to the server.
-The server aggregates these gradients and updates the model parameters using
+They then compute local updates on their respective local datasets, transmitting the local gradients $\nabla f_{j}(\bm{w}^{t})$ back to the server.
+The server aggregates these gradients and updates the model using
 $\bm{w}^{t+1}=\bm{w}^{t}-\eta_{t}\sum_{j\in\mathbb{S}^{t}}\theta^{t}_{j}\nabla f_{j}(\bm{w}^{t}).$ (1)
 Here, $\eta_{t}$ represents the predefined learning rate, and $\theta^{t}_{j}$ denotes the weight assigned to client $j\in\mathbb{S}^{t}$ in the $t$-th round.
-The training process lasts for $T$ communication rounds until the global model achieves convergence.
-Here, the subset $\mathbb{S}^{t}$ and the weight $\theta^{t}_{j}$ with $j\in\mathbb{S}^{t}$ are the variables to be determined.
+The training process lasts for $T$ rounds until the global model achieves convergence.
+The subset $\mathbb{S}^{t}$ and the weights $\theta^{t}_{j}$ with $j\in\mathbb{S}^{t}$ are the variables to be determined.
 III-B The Optimization Function
-III-B1 Client Selection in a Single Round
-In a single round, we identify the subset of clients that best approximates the data distribution of the full set of clients.
+III-B1 Formulation of Principle I
+In our work, we employ the Principle I to guide the client selection in a single round.
 Specifically, we evaluate the estimation error between the aggregated gradient obtained from the client subset $\mathbb{S}^{t}$ and the aggregated gradient obtained from the full client set $\mathbb{N}$.
@@
 In Eq.
-(2), when clients are not selected in the $t$-th round, the gradients $\nabla f_{i}(\bm{w}^{t})$ for these clients become unknown, rendering the optimization of $\text{D}(\mathbb{S}^{t})$ intractable.
-To address this challenge, we obtain the upper bound of $\text{D}(\mathbb{S}^{t})$, denoted as $\text{DUB}(\mathbb{S}^{t})$ in Eq.
-(3), and minimize this upper bound instead.
-The proof of Theorem 1 is provided in the supplementary file.
-Theorem 1.
-Define
-$\text{DUB}(\mathbb{S}^{t})\triangleq\sum_{i=1}^{N}\min_{j\in\mathbb{S}^{t}}\|\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})\|_{2}^{2},$ (3)
-then $\text{DUB}(\mathbb{S}^{t})$ serves as an upper bound for $\text{D}(\mathbb{S}^{t})$.
-Specifically, following the analysis in [32], we assume a mapping $\xi^{t}:\mathbb{N}\rightarrow\mathbb{S}^{t}$ that assigns each client $i\in\mathbb{N}$ to a client $j\in\mathbb{S}^{t}$, i.e., $\xi^{t}(i)=j$, indicating that client $i$ can be approximately represented by client $j$ in the $t$-th round.
+(2), when clients are not selected in the $t$-th round, their gradients $\nabla f_{i}(\bm{w}^{t})$ become unknown.
+To address this challenge, following the analysis in [30], we assume a mapping $\xi^{t}:\mathbb{N}\rightarrow\mathbb{S}^{t}$ that assigns each client $i\in\mathbb{N}$ to a client $j\in\mathbb{S}^{t}$, i.e., $\xi^{t}(i)=j$, indicating that client $i$ can be approximately represented by client $j$ in the $t$-th round.
 For a client $j\in\mathbb{S}^{t}$, let $\mathbb{C}^{t}_{j}=\{i\in\mathbb{N}\,|\,\xi^{t}(i)=j\}$ be the set of clients that can be represented by client $j$.
@@
 In the later section, we utilize $\text{DUB}(\mathbb{S}^{t})$ when minimizing the estimation error $\text{D}(\mathbb{S}^{t})$ is required.
-To manage the computation of $\text{DUB}(\mathbb{S}^{t})$ in Eq.
-(3), we introduce the metric $\text{Dist}_{i,j}(t)$ to quantify the distance between two clients $i$ and $j$ in the $t$-th round.
-This metric is defined as
-$\text{Dist}_{i,j}(t)=||\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})||_{2}^{2}.$ (4)
-Note that when clients $i,j\in\mathbb{N}$ are not selected in the $t$-th round, their gradients $\nabla f_{i}(\bm{w}^{t})$ and $\nabla f_{j}(\bm{w}^{t})$, and consequently $\text{Dist}_{i,j}(t)$, cannot be calculated.
-To address this issue, we propose partial updates for $\text{Dist}_{i,j}(t)$.
+The proof of Theorem 1 is provided in supplementary file.
+Theorem 1.
+Define
+$\text{Dist}_{i,j}(t)=||\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})||_{2}^{2},$ (3)
+and
+$\text{DUB}(\mathbb{S}^{t})$ $\triangleq\sum_{i=1}^{N}\min_{j\in\mathbb{S}^{t}}\text{Dist}_{i,j}(t),$ (4)
+then $\text{DUB}(\mathbb{S}^{t})$ serves as an upper bound for $\text{D}(\mathbb{S}^{t})$.
+In our work, we use partial updates to compute $\text{Dist}_{i,j}(t)$.
 Specifically, when $t=0$, all clients are selected, and we compute $\text{Dist}_{i,j}(t)$ for each pair of clients.
 Then, for $t\geq 1$, the server updates $\text{Dist}_{i,j}(t)$ using the gradients only from the selected $K$ clients.
-That is expressed as
+That is,
 $\text{Dist}_{i,j}(t)=\begin{cases}&\|\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})\|_{2}^{2},\;\text{if}\;i,j\in\mathbb{S}^{t},\\ &\text{Dist}_{i,j}(t-1),\;\text{otherwise}.\end{cases}$ (5)
-Although using $\text{Dist}_{i,j}(t-1)$ to approximate $\text{Dist}_{i,j}(t)$ may introduce biases when either client is not selected, experimental results demonstrate that there are minimal impacts on the model convergence.
-This can be attributed to two reasons.
-First, there are limited variations in gradients between successive communication rounds.
-Moreover, the individual fairness constraint proposed in Section III-B2 ensures that each client has the opportunity of being selected and the corresponding $\text{Dist}_{i,j}(t)$ can be updated.
-In the following, we use $\text{Dist}_{i,j}(t)$ in Eq.
-(5) to replace the term $\|\nabla f_{i}(\bm{w}^{t})-\nabla f_{j}(\bm{w}^{t})\|_{2}^{2}$ in Eq.
-(3) to make $\text{DUB}(\mathbb{S}^{t})$ solvable.
-III-B2 Client Selection Across Multiple Rounds
-In our work, we propose a novel fairness constraint to guide the client selection process across multiple rounds.
-This constraint asserts that clients with similar data distributions should have similar frequencies of being selected.
-To formulate the fairness constraint, we establish metrics to measure the similarity between a pair of clients in terms of their data distributions and the frequencies of being selected, respectively.
+Although using $\text{Dist}_{i,j}(t-1)$ to approximate $\text{Dist}_{i,j}(t)$ may introduce biases, experimental results demonstrate that there are minimal impacts on the model convergence.
+This is because there is limited gradient variations between successive communication rounds, and the proposed individual fairness constraint ensures that all clients have the opportunity of being selected and the corresponding $\text{Dist}_{i,j}(t)$ can be updated.
+III-B2 Formulation of Principle II
+As mentioned in Section I, we propose the individual fairness constraint (Principle II) to guide the client selection across multiple rounds.
+It asserts that clients with similar data distributions should have similar frequencies of being selected.
 First, we use $\text{Dist}_{i,j}(t)$ in Eq.
-(4) to measure the similarity of the data distribution and use Eq.
-(5) to compute its value.
+(3) to measure the similarity of the data distribution.
 Second, let $x_{i,t}$ represent whether the $i$-th client is selected in the $t$-th round, with $x_{i,t}=1$ if $i\in\mathbb{S}^{t}$ and $x_{i,t}=0$ otherwise.
@@
 $p_{i}=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}(x_{i,t}).$ (6)
-to evaluate the frequency of a client $i$ being selected throughout the total $T$ rounds.
-Note that $\{x_{i,t}\}$ forms a stochastic process, and the expectation operation $\mathbb{E}$ is applied in Eq.
-(6).
-Given two clients $i,j\in\mathbb{N}$, we use $|p_{i}-p_{j}|$ to quantify the difference between their frequencies of being selected.
-Building on $\text{Dist}_{i,j}(t)$ and $p_{i}$, we employ the $\epsilon\text{-}\delta$-IF framework utilized in [35, 36] to quantify the fairness constraint.
-We evaluate a client selection strategy at the end of the $T$-th communication round.
-Specifically, given $\epsilon,\delta\geq 0$, a client selection strategy is of individual fairness if for every pair of clients $i,j\in\mathbb{N}$ with $\text{Dist}_{i,j}(T)\leq\epsilon$, the difference in their frequencies of being selected, namely, $p_{i}$ and $p_{j}$, should not exceed $\delta$.
-The mathematical formulation of $\epsilon\text{-}\delta$-IF is provided in Definition 1, and the impact of parameter selection for $\epsilon$ and $\delta$ is discussed in Section V-C.
+to evaluate the frequency of a client $i$ being selected in $T$ rounds. Note that $\{x_{i,t}\}$ forms a stochastic process, and the expectation operation $\mathbb{E}$ is applied in Eq. (6).
+Building on $\text{Dist}_{i,j}(t)$ and $p_{i}$, we employ the $\epsilon\text{-}\delta$-IF framework utilized in [33, 34] to quantify the individual fairness constraint.
+We evaluate a client selection strategy at the end of the $T$-th round.
+Specifically, given $\epsilon,\delta\geq 0$, a client selection strategy is of individual fairness if for any clients $i,j\in\mathbb{N}$ with $\text{Dist}_{i,j}(T)\leq\epsilon$, the difference in their selection frequencies, namely, $p_{i}$ and $p_{j}$, should not exceed $\delta$.
+The mathematical formulation of $\epsilon\text{-}\delta$-IF is provided in Definition 1, and the impact of parameter selection for $\epsilon$ and $\delta$ is discussed in Section V-D.
 Definition 1.
@@
 $\forall\,i,j\in\mathbb{N},\,\,\text{Dist}_{i,j}(T)\leq\epsilon\,\,\Rightarrow\,\,|p_{i}-p_{j}|\leq\delta.$ (7)
-Note that $\epsilon\text{-}\delta$-IF requires examining Eq.
-(7) for all pairs of clients, leading to significant computational overhead and rendering subsequent optimization challenging.
-To address this issue, we propose to determine a reference client $i^{\star}$ for each client $i$, which exhibits the largest difference in the frequency of being selected among clients whose gradient distances are less than $\epsilon$.
+The $\epsilon\text{-}\delta$-IF requires examining Eq.
+(7) for all pairs of clients, leading to significant computational overhead.
+To address this issue, we propose to determine a reference client $i^{\star}$ for each client $i$, which exhibits the largest difference in the selection frequency and has the gradient distance less than $\epsilon$.
 That is,
 $i^{\star}=\text{argmax}_{\text{Dist}_{i,j}(T)\leq\epsilon}\,|p_{i}-p_{j}|,\;\forall\;j\in\mathbb{N}.$ (8)
-Then, given a client $i$ and its reference client $i^{\star}$, a client selection strategy is certified as $\epsilon\text{-}\delta$-IF if
+Then, the $\epsilon\text{-}\delta$-IF is simplified as
 $|p_{i}-p_{i^{\star}}|\leq\delta,\quad\forall\;i\in\mathbb{N}.$ (9)
@@
 In our work, the solution to address this misalignment is proposed in Section III-C.
-III-C Problem Transformation Under Lyapunov Optimization
+III-C Transformation Under Lyapunov Optimization
 To solve P1 in Eq.
 (10), we leverage Lyapunov optimization, a technique from control theory used to analyze the stability of dynamic systems [35, 8].
-The main idea behind Lyapunov optimization is to break down the long-term time-averaged constraints in the optimization problem into constraints that can be adhered to in each communication round.
-This allows us to optimize the objective function directly in each round without violating the constraints specific to that round.
-Furthermore, by ensuring that these constraints are adhered to in each round, we can automatically achieve compliance with the long-term time-averaged constraints at the end of the $T$ round.
+The main idea behind Lyapunov optimization is to break down the long-term time-averaged constraints into constraints that can be adhered to in each communication round.
 By leveraging Lyapunov optimization, the problem stated as P1 in Eq.
 (10) is ultimately converted into the problem P3 in Eq.
-(23), which can be efficiently solved online in each communication round.
-Details of problem transformation are described below.
-Before employing Lyapunov optimization, we remove the absolute value sign from the constraint in Eq. (9) and rephrase it as two equivalent constraints,
+(23).
+Details of the problem transformation are described below, which consists of four steps.
+(a) Transformation of Individual Fairness Constraints.
+Before employing Lyapunov optimization, we remove the absolute value sign from the constraint in Eq.
+(9) and rephrase it as two equivalent constraints,
 $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}(x_{i,t}-x_{i^{\star},t})-\delta\leq 0,\;\forall\,i\in\mathbb{N},$ (11)
 $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}(-x_{i,t}+x_{i^{\star},t})-\delta\leq 0,\;\forall\,i\in\mathbb{N}.$ (12)
-Lyapunov optimization typically involves three steps [37].
-First, a virtual queue is defined for each time-averaged constraint.
-This step converts the time-averaged constraints into stability constraints for virtual queues.
-Second, a Lyapunov function is defined based on virtual queues, and the Lyapunov drift is introduced.
-This step breaks down the time-averaged constraints into requirements for Lyapunov drift in each communication round.
-Finally, the Lyapunov drift is integrated into the optimization objective, allowing us to minimize both the Lyapunov drift and the optimization objective simultaneously.
-In the following, we illustrate how to transform problem P1 using these three steps.
-First, we define two virtual queues, namely $Z_{i}(t)$ for the constraint in Eq.
+(b) Introduction of Virtual Queues $Z_{i}(t)$, $Q_{i}(t)$ and $\Theta(t)$.
+Following the general process of Lyapunov optimization [35], we define two virtual queues, namely $Z_{i}(t)$ for the constraint in Eq.
 (11) and $Q_{i}(t)$ for the constraint in Eq.
@@
 By the introduction of $Z_{i}(t)$ and $Q_{i}(t)$, we have Theorem 2, which converts the long-term constraints on $x_{i,t}$ in Eq.
-(11) and (12) into the stability constraints for $Z_{i}(t)$ and $Q_{i}(t)$.
+(11) and (12) into the stability constraints for $Z_{i}(t)$ and $Q_{i}(t)$, respectively.
 The proof of Theorem 2 is provided in the supplementary file.
@@
 $\lim_{T\to+\infty}\frac{\mathbb{E}[Z_{i}(T)]}{T}=0,\;\text{and}\;\lim_{T\to+\infty}\frac{\mathbb{E}[Q_{i}(T)]}{T}=0.$ (14)
-To make $Z_{i}(t)$ and $Q_{i}(t)$ remain stable, an intuitive approach is to bound the increase of these queues in each round so that they could not grow to infinity.
-Based on this intuition, in our work, we define a global queue $\Theta(t)$, which stores the state of $Z_{i}(t)$ and $Q_{i}(t)$ for all clients.
-That is,
+Then, we define a global queue $\Theta(t)$, which stores the state of $Z_{i}(t)$ and $Q_{i}(t)$ for all clients. That is,
 $\Theta(t)\triangleq\left[Z_{1}(t),\cdots,Z_{n}(t),Q_{1}(t),\cdots,Q_{n}(t)\right].$ (15)
-Second, following Lyapunov optimization, we define the Lyapunov function [37] as
+(c) Introduction of Lyapunov Function $L\left(\Theta(t)\right)$ and Lyapunov Drift $\Delta\left(\Theta(t)\right)$.
+Following the general process of Lyapunov optimization [35], we define the Lyapunov function [35] as
 $L\left(\Theta(t)\right)\triangleq\frac{1}{2}\sum_{i=1}^{N}\left[Z_{i}^{2}(t)+Q_{i}^{2}(t)\right],$ (16)
@@
 (12) into the specific requirements for $\Delta\left(\Theta(t)\right)$ in each communication round.
-At last, to ensure that $\Delta\left(\Theta(t)\right)$ remains sufficiently small, a straightforward method to combine the optimization objection with the drift $\Delta\left(\Theta(t)\right)$, and minimize both simultaneously.
+(d) Introduction of the Tradeoff Factor $V$ and two Variables $m_{i,t}$ and $n_{i,t}$.
+To ensure $\Delta\left(\Theta(t)\right)$ remains sufficiently small, a straightforward method is to combine the optimization objection with the drift $\Delta\left(\Theta(t)\right)$, and minimize both simultaneously.
 Mathematically, in our work, the optimization problem in Eq.
@@
 where $B$ is a constant.
-Based on Theorem 3, the problem P2 in Eq. (18) is transformed into
+The Final Formulation. Based on Theorem 3, the problem P2 in Eq. (18) is transformed into
 $\min_{\mathbb{S}^{t}}\;(1-V)\sum_{i=1}^{N}\left[Z_{i}(t)m_{i,t}+Q_{i}(t)n_{i,t}\right]+V\cdot\text{DUB}(\mathbb{S}^{t}).$ (22)
@@
 […]
-where $Z_{i}(t)$ and $Q_{i}(t)$ are defined in Eq.
-(13), $m_{i,t}$ and $n_{i,t}$ are defined in Eq.
-(20), and $V$ is a hyperparameter that balances the optimization function and the fairness constraint.
-With an appropriate $V$, the solution in Eq.
-(23) can infinitely approximate the optimal solution in Eq.
-(18) [38].
-In the following, we select clients by minimizing $G(\mathbb{S}^{t})$ in Eq.
-(23) in each round.
+which is the final optimization function. In the following, we select clients by minimizing $G(\mathbb{S}^{t})$ in Eq. (23) in each round.
 IV The Proposed LongFed
@@
 To address this issue, we exploit the submodular nature of $G(\mathbb{S}^{t})$.
+Algorithm 1 The Proposed Client Selection Strategy
+1:Input: $Z_{i}(t)$, $Q_{i}(t)$
+2:Output: the selected subset $\mathbb{S}^{t}$
+3:Initialize $\mathbb{S}_{0}^{t}=\emptyset$, $\mathbb{P}_{0}=\mathbb{N}$, and $e=1$.
+4:for  $k\in[0,K-1]$ do
+5:     Determine the reference client $i^{\star}$ for $i\in\mathbb{P}_{k}$
+6:     Compute $m_{i,t}$ and $n_{i,t}$ for $i\in\mathbb{P}_{k}$
+7:     Calculate $G(\mathbb{S}_{k}^{t}\cup\{i\}),\forall\,i\in\mathbb{P}_{k}$
+8:     Identify the client $i_{\text{max}}=\text{argmax}_{i}\overline{G}(\mathbb{S}_{k}^{t}\cup\{i\})$
+9:     $\mathbb{S}_{k+1}^{t}\leftarrow\mathbb{S}_{k}^{t}\cup i_{\text{max}},\;\mathbb{P}_{k+1}\leftarrow\mathbb{P}_{k}\backslash i_{\text{max}}$
+10:end for
 Specifically, a set function $g:2^{\mathbb{N}}\rightarrow\mathbb{R}$ is submodular if for every $A\subseteq B\subseteq\mathbb{N}$ and $\𝐵i\in\mathbb{N}\backslash B$ it holds that $g(A\cup\{i\})-g(A)>g(B\cup\{i\})-g(B)$.
-One typical example of a submodular function is the facility location function [40].
+One typical example of submodular function is the facility location function [37].
 Suppose we aim to select locations from a set of positions =$\mathbb{N}=\{1,\ldots,N\}$ to open facilities and to serve a collection of $K$ users.
@@
 (25).
-Algorithm 1 The Proposed Client Selection Strategy
-1:Input: $Z_{i}(t)$, $Q_{i}(t)$
-2:Output: the selected subset $\mathbb{S}^{t}$
-3:Initialize $\mathbb{S}_{0}^{t}=\emptyset$, $\mathbb{P}_{0}=\mathbb{N}$, and $e=1$.
-4:for  $k\in[0,K-1]$ do
-5:     Determine the reference client $i^{\star}_{t}$ for $i\in\mathbb{P}_{k}$
-6:     Compute $m_{i,t}$ and $n_{i,t}$ for $i\in\mathbb{P}_{k}$
-7:     Calculate $G(\mathbb{S}_{k}^{t}\cup\{i\}),\forall\,i\in\mathbb{P}_{k}$
-8:     Identify the client $i_{\text{max}}=\text{argmax}_{i}\overline{G}(\mathbb{S}_{k}^{t}\cup\{i\})$
-9:     $\mathbb{S}_{k+1}^{t}\leftarrow\mathbb{S}_{k}^{t}\cup i_{\text{max}},\;\mathbb{P}_{k+1}\leftarrow\mathbb{P}_{k}\backslash i_{\text{max}}$
-10:end for
 Prior studies show that the greedy algorithm is an effective solution for finding the maximum value of a submodular function [38, 36].
@@
 IV-B Convergence Analysis
-We provide a theoretical analysis to demonstrate the convergence ability of the proposed client selection strategy.
+To implement the theoretical analysis, we establish six assumptions regarding the local models and data distribution heterogeneity among clients.
 The analysis uses FedAvg as the aggregation method, and it can be extended to other federated optimization methods as well.
-To implement the theoretical analysis, we establish six assumptions regarding the local models and data distribution heterogeneity among clients.
-First, based on the effective greedy selection strategy in Algorithm 1, we assume that the estimation error between the client subset and the full client set (Eq.
+First, we assume that the estimation error between the client subset and the full client set (Eq.
 (2)) is small and can be quantified by a variable $\rho$, as stated in Assumption 1.
-This assumption has also been studied in the context of data coreset selection [32] and client selection methods [27].
 Note that $\rho$ is used as a measure to characterize the quality of the estimation, and the analysis holds for any $\rho<\infty$.
-Assumption 1.
-At each communication round $t$, we assume that the gradient aggregated from the selected subset of clients can provide a good approximation of the gradient aggregated from the full set, i.e.,
+Assumption 1. At a round $t$, we assume that the gradient aggregated from the selected subset of clients can provide a good approximation of the gradient aggregated from the full set, i.e.,
 $\left\|\sum_{i\in\mathbb{N}}\nabla f_{i}(\bm{w}^{t})-\sum_{j\in\mathbb{S}^{t}}\theta_{j}^{t}\nabla f_{j}(\bm{w}^{t})\right\|\leq\rho.$ (26)
@@
 The proof is provided in the supplementary file.
-The convergence rate $\mathcal{O}(1/t)$ is the same as existing client selection methods [27, 18, 43].
+The convergence rate $\mathcal{O}(1/t)$ is the same as loss-guided client selection methods [16, 12, 40].
 Theorem 4.
@@
 $\mathbb{E}\|\bm{w}^{\star}-\bm{w}^{t}\|_{2}^{2}\leq\mathcal{O}(1/t)+\mathcal{O}(\rho).$ (32)
-The term $\rho$ encodes the estimation error in Eq.
-(2).
-When more clients are selected in a communication round, the subset of selected clients can better approximate the data distribution of the full client set, then the term $\rho$ is smaller.
-Particularly, $\rho$ becomes zero when all clients are selected in each round, i.e., $K=N$.
-We also analyze the impact of varying $K$ empirically and propose an adaptive client selection method [44].
-However, to reduce communication overhead, the number of selected clients is usually less than the total number of available clients, i.e., $K<N$.
-In this case, $\rho$ remains a non-vanishing term.
-In experiments, we observe that the proposed client selection strategy guides the system to converge along a trajectory similar to that of full client participation, but there exists a non-diminishing solution bias dependent on $\rho$.
+In Eq.
+(2), the term $\rho$ encodes the estimation error.
+When more clients are selected, the term $\rho$ is smaller.
+Particularly, $\rho$ becomes zero when all clients are selected, i.e., $K=N$.
+In practical settings, as only limited clients can be selected, $\rho$ remains a non-vanishing term.
+In our experiments, we also observe that there exists a non-diminishing solution bias dependent on $\rho$.
 This observation is consistent with our theoretical analysis.
+The impact of varying $K$ is also empirically analyzed in [41].
 V Experiment
 V-A Experimental Settings
-We consider the cross-device federated learning scenario where a large number of clients exist, each with limited computational power.
-In our experiments, we fix the number of clients at $N=100$ and employ FedAvg [1] as the aggregation method.
-The proposed client selection strategy is evaluated on two datasets, FMNIST [45] and CIFAR-10 [46].
-Following the approach in [18], we opt for basic and small-scale models to accommodate the clients’ restricted computational resources.
-Specifically, for the FMNIST dataset, we utilize a multilayer perception (MLP) with two hidden layers as the global model.
-For CIFAR-10, our global model adopts a convolutional neural network (CNN) architecture consisting of three convolutional layers, with the outputs of these layers fed into an MLP layer.
-Further details on training and hyperparameter settings can be found in the supplementary file.
-We explore four data partitioning methods, including an independently and identically distributed (IID) approach and three heterogeneous approaches.
-In the IID approach, we randomly partition the dataset into $N$ parts, with each client assigned one part.
-We select $K=5$ clients in each round.
-Three heterogeneous data partitioning approaches are as follows.
+We consider the cross-device federated learning scenario where $N=100$ clients exist, each with limited computational power.
+The FedAvg method [1] is used as the aggregation method.
+We evaluate our method on two datasets, FMNIST [42] and CIFAR-10 [43].
+Following [12], we opt for basic and small-scale models to accommodate the clients’ restricted computational resources.
+Specifically, for the FMNIST dataset, we utilize a multilayer perception (MLP) with two hidden layers.
+For CIFAR-10, we adopts a convolutional neural network (CNN) architecture consisting of three convolutional layers.
+Details of training and hyperparameter settings can be found in supplementary file.
+Data Partition Methods.
+We explore four data partitioning methods.
+The first is an independently and identically distributed (IID) approach, where we randomly partition the dataset into $N$ parts, with each client assigned one part.
+We also consider three heterogeneous data partitioning approaches as follows.
 (i) 1 Shard Per Client (1SPC).
-Following the setting in [1], we divide the dataset into $N$ shards, ensuring that data within a shard shares the same label.
+Following [1], we divide the dataset into $N$ shards, ensuring that data within a shard shares the same label.
 We randomly assign a shard to each client so that a client has data with only one label.
-This is the most heterogeneous case.
 In this case, we select $K=10$ clients in each round.
 (ii) 2 Shards Per Client (2SPC).
-Similar to the 1SPC setting, the dataset is partitioned into $2N$ shards, where data within each shard shares the same label.
-Clients are randomly assigned two shards, where most clients have data with two labels (two shards with different labels), while a few clients have data with only one label (two shards with the same label).
+The dataset is partitioned into $2N$ shards, with each shard containing data that shares the same label.
+Clients are randomly assigned two shards, allowing them to have data with as most two distinct labels.
 In this case, we set $K=5$ in each round.
@@
 We partition the dataset based on a Dirichlet distribution parameterized by a concentration variable $\alpha$, where a smaller value of $\alpha$ indicates higher heterogeneity.
-In our work, we set $\alpha=0.8$.
-Under this setting, clients have data with all labels, but the data size of each label varies among clients.
-We determine the data size of each label for each client following the analysis in [47, 18].
+In our work, we set $\alpha=0.8$, and determine the data size of each label for each client following [44, 12].
 In this case, we choose $K=5$ clients in each round.
-Figure 2: Test accuracy on FMNIST and CIFAR-10 under the IID scenario.
+We also experimented with smaller values, such as $\alpha=0.5$ and $\alpha=0.1$.
+We found that our method consistently achieved the best model performance.
 Note that the 1SPC and 2SPC scenarios address label shift heterogeneity, while the Dir scenario accounts for heterogeneity in both labels and dataset size.
-Furthermore, apart from the assumption of a fixed number of clients ($K$) to be selected, we also explore the impact of varying $K$ and propose an adaptive method to determine its value.
-Details of the adaptive method can be found in [44].
+Figure 2: Test accuracy under the IID scenario.
+Comparison Methods.
+We first evaluate the effectiveness of the proposed method by comparing it with existing client selection strategies in terms of model performance.
+The baselines are as follows.
+(1) Random selection strategy (Random, 2017) [1]; (2) Loss-guided selection methods, including active client selection strategy (AFL, 2019) [10], power-of-choice selection strategy (Powerd, 2022) [11], and diverse client selection strategy (divFL, 2022) [16]; (3) Contribution-based methods, including the Shapley value-based method (ShapleyFL, 2023) [22] and ranking-based client selection (FedRank, 2024) [45].
+Cluster-based methods are not considered, as they require explicit clustering patterns, which are not applicable in the 2SPC and Dir scenarios.
+Additionally, we include full participation method (Full) with $K=100$ as a reference.
+Furthermore, we evaluate the proposed method in terms of fairness.
+As introduced in Section II, the uniform selection constraint from [29] (2023) in performance fairness research is most aligned with our work.
+In our study, we adopt this uniform selection constraint and apply it to loss-guided selection methods, given their superior model performance.
+Specifically, we apply the uniform selection constraint to AFL [10], Powerd [11], and divFL [16], denoting them as AFL+Fair, Powerd+Fair, and divFL+Fair, respectively.
+We then compare our method against these three baselines in terms of fairness.
 Figure 3: Test accuracy on FMNIST and CIFAR-10 under three heterogeneous data partitioning settings.
-TABLE I: Time overhead analysis on FMNIST in the 1SPC scenario.
-Method Random AFL Powerd DFairFed
-Time (ms) For Client Selection Only 0.038 0.153 0.016 0.529
-Time (ms) For Completing a Communication Round 2019 2044 2033 2472
-We conduct the experiments five times using different randomized initial conditions and present the average results.
-In our experiments, we set $\delta=0.1$ , $\epsilon=0.3$, and $V=0.9$ for FMNIST, and $\delta=0.05$, $\epsilon=0.05$, $V=0.7$ for CIFAR-10.
-We compare the proposed LongFed with three baseline methods: (1) the random selection strategy (Random) [1]; (2) active client selection strategy (AFL), where clients with higher training losses have higher probabilities of being selected [16]; (3) power-of-choice strategy (Powerd), where clients are selected based on the power of $d$ choices load balancing strategy [17].
-These methods are commonly used as baselines in [18, 27].
-Additionally, we compare our method with the full client participation method (Full) with $K=100$.
+Note that the divFL [16] method also selects clients to best represent full client participation, which aligns with our proposed Principle I.
+However, divFL does not consider fairness.
+By comparing divFL with our proposed longFed, we can evaluate the impact of our proposed individual fairness on improving model performance.
+Additionally, by comparing divFL+Fair with longFed, we can further evaluate how our proposed individual fairness outperforms the uniform selection constraint.
 V-B Experimental Results
-V-B1 Convergence Analysis
+V-B1 Model Performance
 The experimental results under the IID scenario are presented in Fig.
-2, and the results under three heterogeneous data partitioning methods are presented in Fig.
+2.
+First, as shown in Fig.
+2 (a) and (b), our proposed longFed outperforms existing methods, particularly on CIFAR-10 dataset, validating its effectiveness in enhancing model performance in the IID scenario.
+Second, longFed consistently outperforms divFL, further demonstrating the effectiveness of our proposed individual fairness in the IID scenario.
+Figure 4: Fairness results on the CIFAR-10 dataset in the 2SPC scenario.
+The experimental results under three heterogeneous scenarios are presented in Fig.
 3.
-In Fig.
-2 and Fig.
-3, the $x$ axis represents the communication round, and the $y$ axis displays the test accuracy of the global model.
-For a clearer comparison, we apply the moving average rule [48] with a sliding window size of 10 to smooth the results.
-Fig.
-2 reveals that in the IID scenario, these methods achieve similar performance.
-For three heterogeneous scenarios on the FMNIST dataset, as shown in Fig.
-3 (a), (b), and (c), respectively, the proposed method achieves performance that is the most similar to that with full client participation.
-Moreover, it exhibits faster convergence speed and superior test accuracy compared to prior works.
-Notably, there is an approximate $20\%$ improvement in the 1SPC scenario and an $8\%$ improvement in the 2SPC and Dir scenarios.
-Similar trends are also observed for the CIFAR-10 dataset.
-Figure 4: Visualization of the selected clients on the FMNIST dataset under the 1SPC scenario.
-(a) is the t-SNE plot of client embeddings, where clients are organized into 10 clusters.
-(b)-(f) display the results of the selected clients using different client selection strategies, with the chosen clients marked by black stars.
-Figure 5: Visualization of the selected clients on the FMNIST dataset under the 2SPC scenario.
-(a) shows the t-SNE plot of client embeddings, where the clustering pattern is not evident.
-In (b)-(f), circles are drawn around the selected clients, indicating that clients covered by a circle can be represented by the corresponding selected client.
-V-B2 Time Overhead Analysis
-We evaluate the time overhead of the proposed client selection strategy, and the results are presented in Table I.
-The time overhead analysis is conducted on the FMNIST dataset under the 1SPC scenario, with similar observations across other scenarios.
-In Table I, the first row focuses on the time solely for client selection within a communication round, while the second row denotes the overall time required for a complete communication round, including client selection, local updates, and global aggregation.
-First, examining the time solely for client selection, our proposed method exhibits only a marginal increase (less than 0.4ms) compared to existing methods.
-This increment is acceptable in most real applications.
-In addition, the time required for client selection (about 0.5ms) is almost negligible compared to the total time required to complete a communication round (about 2000ms).
-This highlights that the majority of time is allocated to local updates within clients.
-Second, examining the time required to complete a round, our proposed strategy imposes only a modest additional computational time (about 400 ms) compared to existing approaches.
-This is primarily because evaluating the distance $\text{Dist}_{i,j}(t)$ in Eq.
-(5) is time-consuming, especially when the dimension of $\nabla f_{i}(\bm{w}^{t})$ is high.
-The time overhead can be reduced by employing a more efficient method for calculating distances in high-dimensional gradient space.
-Nevertheless, the marginal increase in time for the proposed strategy is justified when considering its superior performance in client selection.
-V-B3 Visualization of Client Selection Strategy
-We provide visualizations of the selected clients to offer an interpretable analysis of our proposed strategy.
-As shown in Fig.
-4, in the 1SPC scenario on the FMNIST dataset, baseline methods typically select two or more clients from the same cluster.
-In contrast, the proposed LongFed, as shown in 4 (e), selects one client from each cluster, which effectively approximates the data distribution of the full client set.
-Additionally, as shown in Fig.
-4 (f), it chooses a diverse subset of clients in the multi-round selection, adhering to Principle II.
-The results in the 2SPC scenario are illustrated in Fig.
-5, where the clustering pattern is not evident.
-For baseline methods, the circles around selected clients often overlap and cover only a minority of clients.
-However, in our proposed method, as shown in Fig.
-5 (e), these circles cover the majority of clients.
-This suggests that the selected clients provide a better approximation of the full client set compared to the baseline approaches.
-Moreover, as shown in Fig.
-5 (f), the strategy selects a diverse subset of clients across multiple rounds, satisfying Principle II.
-Visualization results of other scenarios are provided in the supplementary file.
-V-C Individual Fairness Analysis
-We analyze the proposed individual fairness in terms of the trade-off factor $V$, the similarity measure $\epsilon$, and the probability difference measure $\delta$.
-The results are illustrated using the 2SPC scenario on the CIFAR-10 dataset, and similar trends are observed for other scenarios.
-The random selection strategy is chosen as the baseline because it selects clients randomly and without bias.
-The impact of these parameters is assessed through two metrics: test accuracy and the standard deviation of the selection probability $p_{i}$ between clients with similar data distributions, denoted as $\sigma(t)$.
-To calculate $\sigma(t)$, we define $c_{i}(t)$ as the accumulated number of selections for client $i$ during previous $t$ communication rounds.
+First, compared to prior works (except for divFL), the proposed longFed exhibits faster convergence and superior test accuracy.
+Notably, it achieves an approximate $20\%$ improvement in the 1SPC scenario and an $8\%$ improvement in the 2SPC and Dir scenarios on the FMNIST dataset.
+Besides, the proposed method consistently achieves performance comparable to full client participation across all three scenarios and both datasets.
+This validates the effectiveness of our method in enhancing model performance.
+Furthermore, compared to divFL, longFed exhibits similar performance in Fig.
+3 (a), (c), and (d), but achieves significant improvements in the other three scenarios.
+This highlights the effectiveness of our proposed individual fairness in improving model performance.
+Notably, in Fig.
+3 (b) and (f), longFed even surpasses full client participation in some rounds.
+This phenomenon typically occurs in the early stages of training due to optimization dynamics and the stochasticity introduced by client selection.
+As training progresses, these effects gradually diminish.
+Additionally, FedRank performs the worst, as it relies solely on pairwise relationships between clients and lacks a global perspective on data diversity.
+This leads to biased selections-for example, it tends to overlook important yet less frequently ranked clients-ultimately resulting in suboptimal training and poorer convergence.
+V-B2 Fairness
+We evaluate the fairness through the standard deviation in the selection probability, denoted as $\sigma(t)$.
+To calculate $\sigma(t)$, we define $c_{i}(t)$ as the accumulated number of selections for client $i$ during previous $t$ rounds.
 That is, $c_{i}(t)=\sum_{k=1}^{t}x_{i,k}$.
@@
 A smaller $\sigma(t)$ indicates that the client selection strategy aligns more closely with the individual fairness constraint.
+We compare the proposed method with three baselines: AFL [10], Powerd [11], and divFL [16], along with their respective versions incorporating the uniform selection constraint, denoted as AFL+Fair, Powerd+Fair, and divFL+Fair.
+These methods are evaluated based on two key aspects: test accuracy, and the standard deviation $\sigma(t)$.
+The analysis is conducted on both the CIFAR-10 and FMNIST datasets under the 2SPC scenario, with similar trends observed in other settings.
+The results of the CIFAR-10 and FMNIST datasets are illustrated in Fig.
+4 and Fig.
+5, respectively.
+In these figures, the vanilla methods are represented by solid lines, while the methods incorporating the uniform selection constraint are indicated by dashed lines.
+TABLE I: Time overhead analysis on FMNIST in the 1SPC scenario.
+Method Random AFL Powerd ShapleyFL FedRank divFL LongFed
+Time (ms) for client selection 0.038 0.153 0.016 221.023 0.057 0.219 0.529
+Time (ms) for a complete round 2019 2044 2033 3242 2045 2407 2472
+In Fig.
+4, the test accuracy results indicate that the proposed method achieves the best performance.
+However, applying the uniform selection constraint leads to a degradation in model performance for all three baselines.
+On the other hand, from the perspective of fairness, incorporating the uniform selection constraint reduces the standard deviation $\sigma(t)$ for all three baseline methods.
+Notably, divFL+Fair achieves the lowest $\sigma(t)$ among all methods.
+The proposed methods attains a comparable but slightly higher standard deviation than divFL+Fair.
+This phenomenon occurs because the uniform selection constraint enforces equal selection probabilities across all clients, whereas the proposed method ensures similar selection probabilities only for clients with similar data distributions.
+As a result, divFL+Fair exhibits slightly better fairness performance than our method.
+Considering both test accuracy and fairness (as measured by $\sigma(t)$), the proposed individual fairness approach demonstrates a superior balance, effectively improving both model performance and fairness simultaneously compared to the uniform selection constraint.
+Figure 5: Fairness results on the FMNIST dataset in the 2SPC scenario.
+In Fig.
+5, the results for the standard deviation $\sigma(t)$ follow a similar trend to that of the CIFAR-10 dataset.
+That is, introducing the uniform selection constraint improve fairness, and the proposed LongFed achieves fairness comparable to divFL+Fair.
+For test accuracy, we observe that applying the uniform selection constraint leads to a decline in performance for Powerd, while AFL maintains similar performance, and divFL exhibits a slight improvement.
+This may be because these methods typically select the same subset of clients across multiple rounds, potentially leading the system to a suboptimal solution.
+The introduction of the uniform selection constraint forces these methods to select different clients, particularly those that were previously under-selected, helping to escape the suboptimal state.
+Most importantly, the proposed LongFed still achieves the best overall performance, demonstating its effectivess in balancing both fairness and model accuracy.
+V-B3 Time Overhead Analysis
+We evaluate the time overhead of the proposed client selection strategy, and the results are presented in Table I.
+The analysis is conducted on the FMNIST dataset under the 1SPC scenario, with similar trends observed across other scenarios.
+In Table I, the first row highlights the time dedicated solely to client selection, while the second row denotes the overall time required for a complete round, including client selection, local updates, and global aggregation.
+First, examining the time specifically for client selection, our proposed method exhibits only a marginal increase (less than 0.4ms) compared to existing methods.
+This slight increase is primarily due to the computational cost of evaluating the distance $\text{Dist}_{i,j}(t)$ in Eq.
+(3).
+However, this can be mitigated by employing more efficient distance computation techniques in high-dimensional gradient space.
+More importantly, the time required for client selection (approximately 0.5ms) is negligible compared to the total time required to complete a round (approximately 2000ms).
+Therefore, the marginal increase in time for the proposed strategy is justified, given its superior improvements in model performance and fairness.
+Additionally, ShapleyFL [22] incurs significantly higher selection time due to the computationally expensive process of Shapley values, which involves combinatorial computations.
+V-C Visualization of Client Selection Strategy
+Figure 6: Visualization of the selected clients on the FMNIST dataset under the 1SPC scenario.
+Figure 7: Visualization of the selected clients on the FMNIST dataset under the 2SPC scenario.
+We provide visualizations of selection results to offer an interpretable analysis.
+In Fig.
+6, (a) is the t-SNE plot of client embeddings, where clients are organized into 10 clusters.
+(b)-(j) display the results of the selected clients using different client selection strategies, with the chosen clients marked by black stars.
+As shown in Fig.
+6 (b)-(f), baseline methods often select two or more clients from the same cluster.
+In contrast, as shown in Fig.
+6 (i), the proposed LongFed selects one client from each cluster, effectively approximating the data distribution of the full client set.
+Additionally, as shown in Fig.
+6 (g) and (h), divFL tends to select the same set of clients across successive rounds.
+In contrast, as shown in Fig.
+6 (i) and (j), the proposed LongFed selects a more diverse subset of clients across multiple rounds, ensuring fairness in multi-round selection.
+Results in the 2SPC scenario are illustrated in Fig.
+7.
+(a) shows the t-SNE plot of client embeddings, where the clustering pattern is not evident.
+In (b)-(j), circles are drawn around the selected clients, indicating that clients covered by a circle can be represented by the corresponding selected client.
+For baseline methods, as shown in Fig.
+7 (b)-(e), the circles around selected clients often overlap and cover only a minority of clients.
+In our proposed method, as shown in Fig.
+7 (i), these circles cover the majority of clients.
+This suggests that the selected clients provide a better approximation of the full client set compared to the baseline approaches.
+both FedRank and divFL repeatedly select the same subset of clients.
+Notably, this subset remains unchanged after 55 rounds and continues to be selected across nearly 250 rounds.
+As a result, the global model is trained on a limited set of clients, leading to biased predictions for the remaining clients.
+This issue is corroborated by the test accuracy results in Fig.3 (b), where divFL and FedRank exhibit the worst performance.
+Additionally, as shown in Fig.
+5, divFL has the highest standard deviation $\sigma(t)$ among the baselines, further highlighting its fairness limitations.
+In contrast, LongFed maintains a more diverse client selection over multiple rounds, ensuring both fairness and improved generalization.
+These findings are consistent with the observations in the 1SPC scenario.
+Additional visualizations for other scenarios are provided in the supplementary file.
+Moreover, as shown in Fig.
+7 (f)-(h), FedRank and divFL select the same subset of clients.
+Notably, this subset remains unchanged after 55 rounds and continues to be selected across nearly 250 rounds for both FedRank and divFL.
+As a result, the global model is trained only on the 5 selected clients, leading to biased predictions for the remaining clients.
+This issue aligns with the test accuracy results in Fig.3 (b), where divFL and FedRank exhibit the worst performance.
+Additionally, as shown in Fig.
+5, divFL has the highest standard deviation $\sigma(t)$ among the baselines, further highlighting its fairness limitations.
+In contrast, LongFed maintains a more diverse client selection over multiple rounds, ensuring both fairness and improved generalization.
+These findings are consistent with the observations in the 1SPC scenario.
+Additional visualizations for other scenarios are provided in the supplementary file.
+V-D Individual Fairness Analysis
+We analyze the $\epsilon\text{-}\delta$-individual fairness in terms of the trade-off factor $V$, the similarity measure $\epsilon$, and the probability difference measure $\delta$.
+The results are illustrated using the 2SPC scenario on the CIFAR-10 dataset, and similar trends are observed for other scenarios.
+We consider two metrics: test accuracy, and the standard deviation $\sigma(t)$ in Eq.
+(34).
+Figure 8: Impact of the trade-off factor $V$.
 Trade-Off Factor $V$.
-We vary the value of $V$ from $0.0$, $0.4$, $0.6$, and $1.0$, and present the results in Fig.
-6.
-First, we observe that the proposed method with these four values converges faster than the random strategy and achieves better performance.
-Among them, $V=0.6$ yields the best test accuracy.
-This suggests that the individual fairness constraint improves the overall performance.
-Second, the random strategy achieves the lowest standard deviation $\sigma(t)$.
-The proposed method exhibits significantly larger $\sigma(t)$ when $V=1.0$.
-This is because when $V=1.0$, our proposed method prioritizes minimizing the estimation error in Eq.
-(2) and neglects the fairness constraint.
+We vary the value of $V$ from $0.0$ to $1.0$, where a smaller $V$ indicates a higher priority for the individual fairness constraint.
+The results are presented in Fig.
+8.
+First, we observe that the proposed method achieves the highest test accuracy when $V=0.6$ and $V=0.8$.
+Second, in terms of fairness, the proposed method exhibits significantly larger standard deviation $\sigma(t)$ when $V=1.0$.
+This is because when $V=1.0$, the method solely prioritizes minimizing the estimation error in Eq.
+(2) while neglecting the fairness constraint.
 Furthermore, the standard deviation $\sigma(t)$ decreases significantly when $V\leq 0.8$.
-It indicates that the constraint in Eq.
-(7) indeed enhances the system fairness.
+Considering both model performance and fairness, we recommend setting $V=0.8$ as the optimal choice.
+Figure 9: Impact of the similarity measure $\epsilon$.
 Similarity Measure $\epsilon$.
-We vary the value of $\epsilon$ across $0.05$, $0.1$, $0.2$, and $0.3$, and observe the corresponding test accuracy and standard deviation, as presented in Fig.
-7.
-First, the results on test accuracy indicate that the proposed method outperforms the random strategy consistently across these $\epsilon$ values.
-On the other hand, a larger $\epsilon$ in the proposed method corresponds to a smaller standard deviation.
-Specifically, the results are comparable to the random selection strategy when $\epsilon>0.05$, and even surpass it when $\epsilon=0.3$.
-This is attributed to the fact that a larger $\epsilon$ enforces more clients to follow the fairness constraint, thus the standard deviation obtained through the proposed strategy gets closer to the random selection strategy.
-Figure 6: Impact of the trade-off factor $V$.
-Figure 7: Impact of the similarity measure $\epsilon$.
+We vary the value of $\epsilon$ across $0.05$, $0.1$, $0.2$, and $0.3$, and evaluate the corresponding test accuracy and standard deviation, as shown in Fig.
+9.
+First, we observe that test accuracy remains relatively consistent and is not particularly sensitive to the choice of $\epsilon$.
+However, a larger $\epsilon$ results in a smaller standard deviation.
+This is because $\epsilon$ defines the similarity threshold for grouping clients, and a larger value considers more clients as similar, leading to reduced variability in selection.
+Considering both model performance and fairness, we recommend setting $\epsilon=0.3$ as the optimal choice.
+Figure 10: Impact of the probability difference measure $\delta$.
 Probability Difference Measure $\delta$.
-We vary the value of $\delta$ across $0.01$, $0.1$, $0.15$, and $0.2$, and observe the corresponding test accuracy and standard deviation, as presented in Fig.
-8.
-The results indicate that the proposed method consistently performs better than the random selection strategy across these $\delta$ values in terms of test accuracy.
-Moreover, a smaller $\delta$ in the proposed method corresponds to a smaller standard deviation.
-This is because a smaller $\delta$ narrows the range of probability difference between clients with similar distributions, resulting in a smaller standard deviation.
+We vary the value of $\delta$ across $0.01$, $0.1$, $0.15$, and $0.2$, and evaluate the corresponding test accuracy and standard deviation, as shown in Fig.
+10.
+The results indicate that test accuracy remains relatively stable and is not sensitive to the choice of $\delta$.
+However, a smaller $\delta$ leads to a lower standard deviation.
+This is because $\delta$ defines the allowable difference in selection probabilities between clients with similar distributions.
+A smaller $\delta$ enforces a stricter constraint, resulting in a smaller standard deviation.
+Given the balance between model performance and fairness, we recommend setting $\delta=0.01$ as the best choice.
+V-E Summary
+In summary, the proposed LongFed demonstrates faster convergence and superior test accuracy, effectively enhancing both model performance and fairness simultaneously.
+Moreover, LongFed consistently selects representative clients to approximate full participation across multiple rounds, regardless of whether a clustering pattern exists among clients.
 VI Conclusion
-In our work, we introduce a novel long-term client selection approach for federated learning, aiming to emulate the optimization trajectory achieved by full client participation.
-We establish two guiding principles to steer the client selection process effectively.
-We formulate the client selection problem as a long-term optimization task, where the objective is to minimize the estimation error in gradient space between the selected subset of clients and the full client set.
-Besides, we impose the individual fairness constraint to ensure similar selection frequencies among clients with similar data distributions.
-Leveraging Lyapunov optimization techniques from control theory, we transform this problem and propose a greedy algorithm to find solutions efficiently.
-Experimental results demonstrate that our proposed strategy closely approximates the performance achieved under full client participation, with significant improvements over previous methods in terms of both accuracy and fairness.
-Additionally, our method exhibits efficiency by incurring minimal time overhead, which is often negligible when compared to local training duration.
-Figure 8: Impact of the probability difference measure $\delta$.
+In this work, we focus on the client selection problem in federated learning, and propose an effective and fair selection method to improve both model performance and fairness.
+To achieve this, we introduce two guiding principles and formulate the client selection problem as a long-term optimization task.
+Experiments show that our method effectively guides the system to converge along a trajectory similar to that of full client participation.
+Visualization results further illustrate that our approach increases data diversity by selecting clients based on their data distributions, thereby improving both model performance and fairness.